### v1

# Order Is All You Need for Categorical Data Clustering

Abstract.
Categorical data composed of nominal valued attributes are ubiquitous in knowledge discovery and data mining tasks. Due to the lack of well-defined metric space, categorical data distributions are difficult to intuitively understand. Clustering is a popular technique suitable for data analysis. However, the success of clustering often relies on reasonable distance metrics, which happens to be what categorical data naturally lack. Therefore, the cluster analysis of categorical data is considered a critical but challenging problem. This paper introduces the new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding the categorical data clusters. To automatically obtain the orders, we propose a new learning paradigm that allows joint learning of clusters and the orders. It turns out that clustering with order learning achieves superior clustering accuracy, and the learned orders provide intuition for understanding the cluster distribution of categorical data. Extensive experiments with statistical evidence and case studies have verified the effectiveness of the new “order is all you need” insight and the proposed method. The source code of this submission is temporarily opened [^1] https://anonymous.4open.science/r/OCL-demo.
Cluster analysis, Categorical data, Order relationship, Distance learning, Partitional clustering
submissionid: 51ccs: Computing methodologies Cluster analysisccs: Computing methodologies Learning latent representationsccs: Computing methodologies Mixture modeling
1. Introduction
[…]
In the literature, there are two main types of clustering algorithms for categorical data: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia and Cheung, 2018) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004) that adjust the sample-cluster affiliations based on certain cluster quality measures. Their performance is commonly dominated by the metric space represented by the adopted measures. Thus, the focus of subsequent research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Naouali et al., 2020).
[…]
The conventional Hamming distance measure and one-hot encoding strategy are very common in categorical data clustering, but they both simply represent the distance between two attribute values based on whether the two values are identical or not, which is far from informative. Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler and Lindenbaum, 2011), inter-attribute relationship (Ahmad and Dey, 2007)(Le and Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang and Cheung, 2022b). On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed Although all the above advanced techniques have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting different clustering tasks on various categorical datasets.
[…]
It is noteworthy that order is the actual basis for clustering in most cases. For instance, when a $k$-means-type algorithm tries to partition a sample into one of the $k$ clusters, it actually obtains the overall order of the $k$ clusters in terms of admitting the sample. The overall order is formed based on the similarity orders of the clusters to the sample reflected by different attributes. Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of values, which makes it easy to obtain the value order. According to the above analysis, we also note that although most existing categorical data clustering approaches do not specify the order of attribute values, their clustering process still passively forms the order through similarity measurement, regardless of whether there is an obvious semantic order of the attribute values.
Therefore, knowing the correct order of attribute values may be beneficial to categorical data clustering, and the experimental results in Figure 1 confirmed this. It can be seen that even randomly generated orders can easily achieve significant accuracy improvements. Another interesting observation is that, although attribute values of some datasets (i.e., the right eight datasets) are with obvious semantic orders, they are rarely optimal for clustering. This is because the semantic orders of attributes are relatively independent of specific clustering tasks, and this information gap can prevent some semantic orders from well-serving clustering. Therefore, obtaining the optimal orders w.r.t. certain clustering tasks is the key to achieving accurate categorical data clustering.
This paper, therefore, proposes a new clustering paradigm that dynamically learns orders based on the current data partition during clustering. Since the order learning is connected to the clustering task, the optimal order for clustering can thus be automatically obtained. To make full use of the learned orders, a distance learning mechanism is also introduced to dynamically measure the sample-cluster affiliation under the guidance of the orders. It turns out that the learned orders can significantly improve clustering accuracy and can also intuitively help understand the distribution and clustering effect of categorical data. Extensive experimental evaluations demonstrate the superiority of the proposed method. The main contributions can be summarized into three points:
  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed for the first time. Accordingly, a new clustering paradigm with order distance learning is proposed.
  • We propose a distance structure learning mechanism to automatically obtain the order and distance. This learning process is jointly performed with clustering, making the clustering performance more robust to various datasets.
  • Both the learning process and the learned order are highly interpretable, which is helpful for understanding complex categorical data. Moreover, the algorithm is efficient, strictly convergent, and can be easily extended to mixed data [^2] Dataset composed of a mixture of categorical and numerical attributes..
2. Related Work
This section overviews the related works in data encoding, representation learning, and consideration of value order, for categorical data clustering.
The data encoding techniques convert the categorical data into numerical values based on certain strategies, and then perform clustering on the numerical encoding. The conventional one-hot encoding and Hamming distance construct similar distance structures for categorical data by assigning distances “1” and “0” to any pair of different and identical values, respectively. As they cannot distinguish different dissimilarity degrees (dos Santos and Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable during clustering. Some statistic-based distance measures (Lin, 1998)(Cheung and Jia, 2013)(dos Santos and Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples. They take a common principle that two similar values may have more consistent statistics within the corresponding attribute or cluster. Some metrics (Ahmad and Dey, 2007)(Le and Ho, 2005) further define distance considering the relevance between attributes. The representative measure proposed in (Ienco et al., 2012) selects more relevant attributes for the distance measurement. Later, the measures (Jia et al., 2016)(Jian et al., 2018a) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have been presented. Coupling-based encoding strategies (Jian et al., 2018c)(Jian et al., 2017) have also been designed accordingly to encode the couplings. All the above techniques perform value-level encoding to reflect sample-level distances, while the work (Qian et al., 2015) directly encodes a data sample by concatenating the distances from it to all the other samples.
Recently, more advanced representation learning methods have been proposed to learn the categorical data encoding. The work proposed in (Zhu et al., 2022) projects dataset through different kernels for more informative coupling representation, and then learns the kernel parameters to obtain task-oriented representation. Some approaches (Jian et al., 2018b)(Zhu et al., 2020) have also been designed to simultaneously encode the numerical and categorical attributes. However, the involved hyper-parameters have a great impact on their performance, and the appropriate parameter selection will be laborious in different real applications. Although a parameter-free learning approach (Jia and Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
Some new advances proposed with the consideration of value order further distinguish categorical attributes into nominal and ordinal attributes. The distance metrics (Zhang et al., 2020)(Zhang and Cheung, 2022b) define the distance between attribute values based on the data statistical information with considering the semantic order. Later, a representation learning approach was introduced by (Zhang and Cheung, 2020), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts. Later, the works (Zhang and Cheung, 2022a)(Zhang and Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering. The most recent work (Zhang et al., 2022) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering. The above recent advances achieve considerable clustering performance improvements and empirically prove the importance of value order in categorical data clustering. However, they all assume the availability of semantic order of attribute values. Moreover, they directly adopt this order for clustering, and overlook that the semantic order may not be suitable under different clustering scenarios as shown in Figure 1.
3. Proposed Method
This section first formulates the problem, then proposes the order distance definition and order learning mechanism, and finally introduces the joint learning of the orders and clusters. Analyses of time complexity and convergence are also provided.
3.1. Problem Formulation
The problem of categorical data clustering with order learning is formulated as follows. Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples (e.g., a collection of $n$ bank clients). Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$ (e.g., occupation, gender, priority, etc.). Each attribute $\mathbf{a}_{r}$ (e.g., occupation) can be denoted as a column vector $\mathbf{a}_{r}=[x_{1,r},x_{2,r},...,x_{n,r}]$ composed of the occupation description of all the $n$ clients. For the categorical attribute “occupation” $\mathbf{a}_{r}$, all its $n$ values are taken from a limited number of qualitative possible values $V_{r}=\{v_{r,1},v_{r,2},...,v_{r,l_{r}}\}$, e.g., $\{\text{lawyer},\text{doctor},...,\text{scientist}\}$. The subscripts indicate the sequential number of an attribute and the sequential number of the attribute’s possible value. For example, doctor is the second value in the $r$-th attribute “occupation”, so we have $v_{r,2}=\text{doctor}$. We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data. Please note that the sequential numbers of possible values only distinguish them, and do not indicate their order.
Assume that there is an optimal order indicating an appropriate underlying distance structure of possible values w.r.t. a clustering task. We denote the optimal order of an attribute $\mathbf{a}_{r}$ as $O^{*}_{r}=\{o^{*}(v_{r,1}),o^{*}(v_{r,2}),...,o^{*}(v_{r,l_{r}})\}$ where the superscript “*” marks the optimum and the value of $o^{*}(v_{r,g})$ is an integer reflecting the unique ranking of $v_{r,g}$ among all the $l_{r}$ possible values in $V_{r}$. By still taking the “occupation” attribute as an example, if its three possible values $v_{r,1}=\text{lawyer}$, $v_{r,2}=\text{doctor}$, and $v_{r,3}=\text{scientist}$, ranking second (i.e., $o(\text{lawyer})=2$), third (i.e., $o(\text{doctor})=3$), and first (i.e., $o(\text{scientist})=1$), respectively. Then the corresponding order of these values are $O_{r}=\{o(\text{lawyer}),o(\text{doctor}),o(\text{scientist})\}=\{2,3,1\}$. Such order reflects a distance structure that roughly satisfies $\text{dist}(v_{r,g},v_{r,h})\propto|o(v_{r,g})-o(v_{r,h})|$. More specifically, the following distance relationship should hold:
  1. (1)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{doctor},\text{lawyer})$,
  2. (2)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{lawyer},\text{scientist})$,
because “doctor” and “scientist” are further away in order, compared to “lawyer” and “doctor” that are adjacent in order.
We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$. Then the order information reflecting the distance structure can guide the partition of data objects during clustering. Our approach, i.e., learning order relations, is more generative than most existing approaches that directly define the distance between attribute values. We focus more on how the intrinsic relationship between values affects the distance between samples in forming clusters, rather than simply defining the clustering-irrelevant distance between values based on data statistics, e.g., occurrence probability of values in an attribute.
For a crisp partitional clustering task to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
(1) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$
where $\mathbf{Q}$ is an $n\times k$ matrix with its $(i,m)$-th entry $q_{i,m}$ indicating the affiliation between sample $\mathbf{x}_{i}$ and cluster $C_{m}$. Specifically, $q_{i,m}=1$ indicates that $\mathbf{x}_{i}$ belongs to $C_{m}$, while $q_{i,m}=0$ indicates that $\mathbf{x}_{i}$ belongs to a cluster other than $C_{m}$, which can be written as:
(2) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$
with $\sum_{i=m}^{k}q_{i,m}=1$. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$. Therefore, the problem to be solved can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$. When the algorithm converges, a sub-optimal solution with $k$ compact bank client clusters is expected to be obtained. In general, “compact” means that the bank clients with more similar characteristics (i.e., with more similar value descriptions across all the attributes) are gathered into the same cluster, and different clusters are relatively distinguishable.
Frequently used symbols in this paper and the corresponding explanations are sorted out in Table 1 below. Then we discuss how to define the order distance and learn order for accurate categorical data clustering.
Table 1. Frequently used symbols. Note that we uniformly use lowercase, uppercase, bold lowercase, and bold uppercase to indicate value, set, vector, and matrix, respectively.
Symbol Explanation
$X$, $A$, $O$, and $C$ Dataset, attribute set, order set, and cluster set
$\mathbf{x}_{i}$ and $x_{i,r}$ $i$-th data object and $r$-th value of $\mathbf{x}_{i}$
$\mathbf{a}_{r}$, $V_{r}$, and $v_{r,g}$ $r$-th attribute, its value set, and $g$-th value of $V_{r}$
$l_{r}$ Total number of possible values in $V_{r}$
$O_{r}$ Order values (integers) corresponding to $V_{r}$
$o(v_{r,g})$ Order value (integer) of possible value $v_{r,g}$
$\mathbf{Q}$ and $C_{m}$ Object-cluster affiliation matrix and $m$-th cluster
$q_{i,m}$ $(i,m)$-th entry of $\mathbf{Q}$ indicating affiliation of $\mathbf{x}_{i}$ to $C_{m}$
$\Theta(\mathbf{x}_{i},C_{m};O)$ Overall order distance between $\mathbf{x}_{i}$ and $C_{m}$
$\theta(x_{i,r},\textbf{p}_{m,r})$ Order distance between $\mathbf{x}_{i}$ and $C_{m}$ reflected by $\mathbf{a}_{r}$
$\mathbf{D}$ and $\mathbf{d}_{i,r}$ Order distance matrix and its $(i,r)$-th entry
$d_{i,r,g}$ Order distance between $x_{i,r}$ and $v_{r,g}$ ($g$-th value of $\mathbf{d}_{i,r}$)
$\mathbf{P}$ and $\mathbf{p}_{m,r}$ Conditional probability matrix and its $(m,r)$-th entry
$p_{m,r,g}$ Occurrence probability of $v_{r,g}$ in $C_{m}$ ($g$-th value of $\mathbf{p}_{m,r}$)
$O_{r,m}$ Order values of $O_{r}$ obtained through $C_{m}$
$L_{r,m}$ A fraction of $L$ jointly contributed by $\mathbf{a}_{r}$ and $C_{m}$
3.2. Order Distance
We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ is defined as:
(3) $\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$
where
(4) $\theta(x_{i,r},\textbf{p}_{m,r})=\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}.$
is the order distance between $\mathbf{x}_{i}$ and all the samples in $C_{m}$ reflected by the $r$-th attribute. $\mathbf{d}_{i,r}$ is an $l_{r}$-dimensional vector containing the order distances between $x_{i,r}$ and possible values in $V_{r}$, and $\mathbf{p}_{m,r}$ is an $l_{r}$-dimensional vector describing the probability distribution of $V_{r}$ within cluster $C_{m}$, which can be written as:
(5) $\mathbf{d}_{i,r}=[d_{i,r,1},d_{i,r,2},...,d_{i,r,l_{r}}]$
and
(6) $\mathbf{p}_{m,r}=[p_{m,r,1},p_{m,r,2},...,p_{m,r,l_{r}}],$
respectively. The $g$-th value of $\mathbf{d}_{i,r}$ is defined as the normalized order distance between $x_{i,r}$ and $v_{r,g}$, which can be written as:
(7) $d_{i,r,g}=\frac{|o(x_{i,r})-o(v_{r,g})|}{l_{r}-1}$
where $l_{r}-1$ in the denominator is to normalize the order distance into the interval $[0,1]$ to ensure distance comparability of different attributes. The $g$-th value of $\mathbf{p}_{m,r}$ is the occurrence probability of $v_{r,g}$ in cluster $C_{m}$, which can be written as:
(8) $p_{m,r,g}=\frac{card(X_{r,g}\cap C_{m})}{card(C_{m})}$
where $X_{r,g}=\{\mathbf{x}_{i}|x_{i,r}=v_{r,g}\}$ is a sample set collecting all the samples in $X$ with their $r$-th values equal to $v_{r,g}$, and the cardinality function $card(\cdot)$ counts the number of elements in a set.
Remark 1.
Rationality of the order distance. The intra-cluster probability distribution and the value order are introduced to form the order distance. The probability descriptor $\mathbf{p}_{m,r}$ is adopted to informatively represent the value distribution in different clusters for appropriate distance measurement and accurate order learning in Section 3.3. By making the distance $\mathbf{d}_{i,r}$ dependent on the order $O$ through Eq. (7), we facilitate a direct connection between the order learning and the clustering tasks. Such a connection leverages the optimization of the order $O$ for more accurate clustering in Section 3.4.
3.3. Order Learning
Based on the above-defined order distance, we discuss how to obtain the optimal order:
(9) $O^{*}=\arg\min_{O}L(\mathbf{Q},O).$
A simple but extremely laborious solution is to search different order combinations across all the attributes. There are a total of $\prod_{r=1}^{d}l_{r}!$ combinations, and each combination should be evaluated on $n$ samples by accumulating their sample-cluster distances calculated by Eq. (3), which makes the search infeasible. A relaxed efficient solution is to search the optimal order within clusters by assuming independence of attributes. Accordingly, optimal order of $\mathbf{a}_{r}$ w.r.t. $C_{m}$ denoted as $O^{*}_{r,m}=\{o^{*}_{m}(v_{r,1}),o^{*}_{m}(v_{r,2}),...,o^{*}_{m}(v_{r,l_{r}})\}$ can be obtained through the efficient intra-cluster traversal search:
(10) $O^{*}_{r,m}=\arg\min_{O_{r}}L_{r,m}$
where $L_{r,m}$ is an inner sum component of the objective $L$. More specifically, $L_{r,m}$ is jointly contributed by the $r$-th attribute and $m$-th cluster, which can be derived from Eqs. (1) and (3) as:
(11) $L_{r,m}=\sum_{\mathbf{x}_{i}\in C_{m}}\theta(x_{i,r},\textbf{p}_{m,r}).$
By searching for the optimal ranking in each cluster and then performing a linear combination based on the cluster size, the overall ranking of each possible value $v_{r,g}$ can be obtained by:
(12) $\bar{o}(v_{r,g})=\sum_{m=1}^{k}\left(o^{*}_{m}(v_{r,g})\cdot\frac{card(C_{m})}{n}\right).$
Since the value of $\bar{o}(v_{r,g})$ is not necessarily an integer, we sort the obtained rankings $\{\bar{o}(v_{r,1}),\bar{o}(v_{r,2}),...,\bar{o}(v_{r,l_{r}})\}$ of an attribute in ascending order to obtain the final ranking of values in $V_{r}$, and the corresponding integer order $\tilde{O}_{r}=\{\tilde{o}(v_{r,1}),\tilde{o}(v_{r,2}),...,\tilde{o}(v_{r,l_{r}})\}$ is obtained. Correspondingly, we collectively refer to all orders obtained through Eqs. (10) - (12) as $\tilde{O}=\{\tilde{O}_{1},\tilde{O}_{2},...,\tilde{O}_{d}\}$.
Remark 2.
Principle of the order learning. In the process of obtaining $\tilde{O}$, the optimal order of each attribute is determined for each cluster by Eqs. (10) - (12). Through Eq. (11), the specific distribution of samples within each cluster is fully considered (see Remark 1) to appropriately reflect the component $L_{r,m}$ of the objective $L$. Hence, an expected minimization on a series of $L_{r,m}$s can be achieved by the determined orders. The overall order of an attribute is obtained by integrating the contributions of all the clusters through Eq. (12). Here, the cluster sizes are taken into account to accommodate more to larger clusters, thereby achieving a more effective reduction on $L$.
3.4. Joint Order and Cluster Learning
We facilitate joint learning of clusters and orders by integrating the order learning into the clustering objective, which can be rewritten in a more detailed form based on Eqs. (1), (3), and (4):
(13) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}$
where the order distance $\mathbf{D}=\{\mathbf{d}_{i,r}|i\in\{1,2,...,n\},r\in\{1,2,...,d\}\}$ is dependent to the order $O$ according to Eqs. (5) and (7), and the probability distribution $\mathbf{P}=\{\mathbf{p}_{m,r}|m=\{1,2...,k\},r\in\{1,2,...,d\}\}$ is dependent to the partition $\mathbf{Q}$ according to Eqs. (6) and (8). Typically, $\mathbf{Q}$ and $O$ are iteratively computed to minimize $L$ following a $k$-modes-type optimization process. Specifically, given $\hat{O}$, the order distance $\mathbf{D}$ can be updated accordingly, and then a new $\mathbf{Q}$ can be computed by:
(14) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{y,r}\\ 0,&\text{otherwise}\\ \end{array}\right.$
with $i=\{1,2,...,n\}$ and $m=\{1,2,...,k\}$. Eq. (14) is a detailed version of Eq. (2) that adopts the order distance defined by Eqs. (3) and (4). Then, $\mathbf{P}$ is updated according to the new $\mathbf{Q}$. When the iterative updating of $\mathbf{Q}$ and $\mathbf{P}$ converges, we obtain $O$ according to Eqs. (10) - (12), and update $\mathbf{D}$ accordingly. In summary, $L$ is minimized by iteratively solving the following two sub-problems:
  1. (1)
    Fix $\hat{\mathbf{Q}}$ and $\hat{\mathbf{P}}$, solve the minimization problem $L(\hat{\mathbf{Q}},O)$ by searching $O$ and updating $\mathbf{D}$ accordingly;
  2. (2)
    Fix $\hat{O}$ and $\hat{\mathbf{D}}$, solve the minimization problem $L(\mathbf{Q},\hat{O})$ by iteratively computing $\mathbf{Q}$, and updating $\mathbf{P}$ according to $\mathbf{Q}$, until convergence;
The whole algorithm named Order and Cluster Learning (OCL) is summarized as Algorithm 1. The next subsection provides analyses of time complexity and convergence to further explain OCL.
Algorithm 1 OCL: Order and Cluster Learning Algorithm
0:  Dataset $X$, number of sought clusters $k$
0:  Partition $\mathbf{Q}$, order $O$
1:  Initialization: Set outer loop counter and convergence mark by $E\leftarrow 0$ and $Conv\_O\leftarrow False$, respectively; Randomly partition $X$ into $k$ clusters $C$, obtain $\mathbf{Q}^{\{I\}}$ and $\mathbf{P}^{\{I\}}$ accordingly; Randomly initialize $O^{\{E\}}$, update $\mathbf{D}^{\{E\}}$ accordingly;
2:  while $Conv\_O=False$ do
3:     $E\leftarrow E+1$; Compute $O^{\{E\}}$ by Eqs. (10) - (12); Update $\mathbf{D}^{\{E\}}$ accordingly; Set inner loop counter and convergence mark by $I\leftarrow 0$ and $Conv\_I\leftarrow False$, respectively;
4:     while $Conv\_I=False$ do
5:        $I\leftarrow I+1$; Compute $\mathbf{Q}^{\{I\}}$ by Eq. (14); Update $\mathbf{P}^{\{I\}}$ accordingly; Obtain $L(\mathbf{Q}^{\{I\}},O^{\{E\}})$;
6:        if $L(\mathbf{Q}^{\{I\}},O^{\{E\}})\geq L(\mathbf{Q}^{\{I-1\}},O^{\{E\}})$ then
7:           $Conv\_I\leftarrow True$; $\mathbf{Q}^{\{E\}}=\mathbf{Q}^{\{I\}}$;
8:        end if
9:     end while
10:     if $L(\mathbf{Q}^{\{E\}},O^{\{E\}})\geq L(\mathbf{Q}^{\{E-1\}},O^{\{E-1\}})$ then
11:        $Conv\_O\leftarrow True$;
12:     end if
13:  end while
3.5. Complexity and Convergence Analysis
We analyze the time complexity to show the scalability of OCL, and then provide its convergence analysis.
Theorem 1.
The time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$.
Proof. Assume it involves $I$ inner iterations (lines 4 - 9 in Algorithm 1) to compute $\mathbf{Q}$ and $\mathbf{P}$, and $E$ outer iterations (lines 2 - 13 in Algorithm 1) to search $O$ and update $\mathbf{D}$. For worst-case analysis, we adopt $\varsigma$ to indicate the maximum number of possible values of a dataset by $\varsigma=\max(l_{1},l_{2},...,l_{d})$.
For each of the $I$ inner iterations, we obtain $\mathbf{P}$ upon the $n$ data samples with complexity $O(nkd\varsigma)$, and obtain $\mathbf{D}$ upon the $\varsigma$ values of $d$ attributes on $n$ samples with complexity $O(nd\varsigma)$. Then the $n$ samples are partitioned to $k$ clusters with complexity $O(nkd\varsigma)$. For $I$ iterations in total, the complexity is $O(Inkd\varsigma)$.
For each of the $E$ outer iterations, a new order of $\varsigma$ possible values of each attribute within each cluster is searched by Eqs. (10) - (12), and the complexity for $d$ attributes and $k$ clusters is thus $O(dkn\varsigma!)$. For $E$ outer iterations with considering the $I$ inner iterations, the overall time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$. □
Remark 3.
Scalability of OCL. The essential characteristic of categorical data is that the number of possible values for an attribute is finite and small, usually ranging from 2 to 8 for real benchmark datasets as shown in the Appendix. Accordingly, $\varsigma$ can be treated as a small constant, and the time complexity is simplified to $O(EInkd)$, which is linear to the scaling of $n$ or $d$.
Theorem 2.
OCL algorithm converges to a local minimum in a finite number of iterations.
Proof. We prove the convergence of the inner and outer loop, i.e., lines 4 - 9 and lines 2 - 13 of Algorithm 1, respectively.
Lemma 0.
The inner loop converges to a local minimum of $L(\mathbf{Q},\hat{O})$ in a finite number of iterations with fixed $\hat{O}$. Note that the number of possible partitions $\mathbf{Q}$ is finite for a dataset with a finite number of objects $n$. We then illustrate that each possible partition $\mathbf{Q}$ occurs at most once by OCL. Assume that two optimizers satisfy $\mathbf{Q}^{\{I_{1}\}}=\mathbf{Q}^{\{I_{2}\}}$, where $I_{1}\neq I_{2}$. Then it is clear that $L(\mathbf{Q}^{\{I_{1}\}},O^{\{E\}})=L(\mathbf{Q}^{\{I_{2}\}},O^{\{E\}})$. However, the value sequence of the objective function $\{L(\mathbf{Q}^{\{1\}},O^{\{E\}}),L(\mathbf{Q}^{\{2\}},O^{\{E\}}),...,L(\mathbf{Q}^{\{I\}},O^{\{E\}})\}$ generated by OCL is strictly decreasing. Hence, the result follows.
Lemma 0.
The outer loop converges to a local minimum of $L(\mathbf{Q},O)$ in a finite number of iterations. We note that the status of orders $O$ is in a discrete space and thus the number of possible orders is finite. The number of possible partitions $\mathbf{Q}$ is also finite. Therefore, the number of possible combinations of $\mathbf{Q}$ and $O$ is finite. We then prove that each possible combination of $\mathbf{Q}$ and $O$ appears at most once by OCL. We note that given $\mathbf{Q}^{\{E_{1}\}}=\mathbf{Q}^{\{E_{2}\}}$, where $E_{1}\neq E_{2}$, the corresponding $O^{\{E_{1}+1\}}$ and $O^{\{E_{2}+1\}}$ can be computed accordingly, satisfying $O^{\{E_{1}+1\}}=O^{\{E_{2}+1\}}$. Then we can obtain their corresponding partitions $\mathbf{Q}^{\{E_{1}+1\}}$ and $\mathbf{Q}^{\{E_{2}+1\}}$ at the convergence of the next inner loop, satisfying $\mathbf{Q}^{\{E_{1}+1\}}=\mathbf{Q}^{\{E_{2}+1\}}$. It is clear that $L(\mathbf{Q}^{\{E_{1}+1\}},O^{\{E_{1}+1\}})=L(\mathbf{Q}^{\{E_{2}+1\}},O^{\{E_{2}+1\}})$. However, the sequence $\{L(\mathbf{Q}^{\{1\}},O^{\{1\}}),L(\mathbf{Q}^{\{2\}},O^{\{2\}}),...,L(\mathbf{Q}^{\{E\}},O^{\{E\}})\}$ is monotonically decreasing. Hence, the result follows.
Since both the inner loop and outer loop of Algorithm 1 converge to a local minimum, convergence of OCL can be confirmed. □
4. Experiments
Different experiments have been designed to evaluate the performance of the proposed order learning strategy as well as the corresponding clustering algorithm. Detailed experimental settings and the obtained results are introduced as follows.
Table 2. Statistics of the 12 datasets. $d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, and $n$ are the numbers of nominal attributes, ordinal attributes, and samples, respectively. $k^{*}$ is the number of true clusters.
No. Dataset Abbrev. $d^{\text{\{nom\}}}$ $d^{\text{\{ord\}}}$ $n$ $k^{*}$
1 Soybean (small) SB 35 0 47 4
2 Nursery School NS 1 7 12960 4
3 Amphibians AP 6 6 189 2
4 Inflammations Diagnosis DS 0 5 120 2
5 Caesarian Section CS 1 3 80 2
6 Hayes-Roth HR 2 2 132 3
7 Zoo ZO 15 0 101 7
8 Breast Cancer BC 5 4 286 2
9 Lymphography LG 15 3 148 4
10 Tic-Tac-Toe TT 9 0 958 2
11 Australia Credit AC 0 8 690 2
12 Congressional Voting VT 16 0 435 2
4.1. Experimental Settings
Four types of experiments have been conducted to evaluate the proposed method. A brief outline of the experiments is given below: (1) Clustering performance evaluation with significance test (Section 4.2): We compared the proposed OCL with conventional and state-of-the-art methods on different types of datasets. To statistically analyze the performance differences between variant approaches, we also conducted a significance test on the clustering results; (2) Ablation studies (Section 4.3): To more specifically demonstrate the effectiveness of the core components of OCL and investigate the role of learned order in improving clustering performance, we conduct ablation studies from the algorithm aspect and data aspect. (3) Convergence and efficiency evaluation (Section 4.4): The values of the objective function after each learning iteration have been recorded to study the convergence of OCL. Its efficiency has also been studied with experiments on large datasets with scaling sizes of samples $n$ and attributes $d$. (4) Visualization of clustering effect (Section 4.5): To visually demonstrate the effectiveness of the order distance metric learned by OCL, we use it and the counterparts to numerically encode the data, and utilize the t-SNE for visualization.
Nine counterparts have been selected for the comparative study. The counterparts include traditional K-MoDes (KMD) algorithm (Huang, 1998), popular entropy-based metric LSM (Lin, 1998), and context-based distance measurement CBDM (Ienco et al., 2009). The latter two are representative measures proposed for categorical data, both of which can be embedded into KMD to conduct clustering. Besides, some state-of-the-art methods have also been taken into account, including Jia’s Distance Metric (JDM) (Jia et al., 2016), Unified Distance Metric based Clustering (UDMC) (Zhang and Cheung, 2022b), HD based Clustering (HDC) (Zhang and Cheung, 2022a), projection based clustering Het2Hom (H2H) (Zhang et al., 2022), Distance Learning based Clustering (DLC) algorithm (Zhang and Cheung, 2020), and graph metric based clustering algorithm ADC (Zhang and Cheung, 2023). Please note that we have set the parameters (if any) of these methods to the values recommended in the corresponding paper. In the experiments, each clustering result was recorded based on the average performance of 10 runs.
Three validity indices, i.e., clustering accuracy (CA) (He et al., 2005), ARI coefficients (Rand, 1971)(Gates and Ahn, 2017), and standardized mutual information NMI (Estévez et al., 2009), have been selected to evaluate the clustering performance. Usually, larger values of these indices indicate better clustering performance. We also conduct the Bonferroni Dunn (BD) test [^3] The Friedman test (Friedman, 1937) is performed first to ensure a meaningful BD post-hoc test. (Demšar, 2006) on the performance of different methods and calculate the critical difference (CD) interval to statistically analyze the superiority of our OCL.
[…]
Table 3. Clustering performance evaluated by CA. “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.8468$\pm$0.15 0.8106$\pm$0.16 0.7894$\pm$0.14 0.7957$\pm$0.15 0.7915$\pm$0.18 0.8000$\pm$0.14 0.9234$\pm$0.12 0.7915$\pm$0.15 0.7660$\pm$0.13 0.9830$\pm$0.05
NS 0.3208$\pm$0.05 0.2975$\pm$0.04 0.2980$\pm$0.04 - 0.3074$\pm$0.05 0.3386$\pm$0.07 0.3113$\pm$0.05 0.3208$\pm$0.05 0.3208$\pm$0.05 0.3573$\pm$0.05
AP 0.5392$\pm$0.03 0.5656$\pm$0.04 0.6159$\pm$0.04 0.5926$\pm$0.03 0.5698$\pm$0.03 0.6032$\pm$0.00 0.5603$\pm$0.03 0.5608$\pm$0.04 0.5529$\pm$0.04 0.6222$\pm$0.00
DS 0.6525$\pm$0.08 0.6525$\pm$0.08 0.7208$\pm$0.12 0.7208$\pm$0.12 0.6867$\pm$0.12 0.6650$\pm$0.12 0.6175$\pm$0.04 0.6867$\pm$0.12 0.6950$\pm$0.12 0.7458$\pm$0.08
CS 0.5462$\pm$0.03 0.5488$\pm$0.06 0.5788$\pm$0.07 0.5875$\pm$0.04 0.5825$\pm$0.05 0.6350$\pm$0.03 0.6075$\pm$0.02 0.5837$\pm$0.05 0.5525$\pm$0.04 0.6313$\pm$0.04
HR 0.3758$\pm$0.02 0.3886$\pm$0.03 0.3917$\pm$0.02 0.3924$\pm$0.04 0.3773$\pm$0.02 0.3697$\pm$0.04 0.3333$\pm$0.00 0.4000$\pm$0.03 0.3758$\pm$0.02 0.4326$\pm$0.02
ZO 0.6624$\pm$0.12 0.6733$\pm$0.11 0.6743$\pm$0.11 0.6752$\pm$0.10 0.6693$\pm$0.10 0.6802$\pm$0.08 0.7059$\pm$0.05 0.6733$\pm$0.10 0.6743$\pm$0.11 0.7792$\pm$0.10
BC 0.5472$\pm$0.04 0.5594$\pm$0.07 0.6392$\pm$0.09 0.6346$\pm$0.09 0.5413$\pm$0.06 0.6140$\pm$0.09 0.5266$\pm$0.06 0.6052$\pm$0.10 0.6185$\pm$0.10 0.6650$\pm$0.05
LG 0.4277$\pm$0.06 0.4446$\pm$0.04 0.4547$\pm$0.06 0.4905$\pm$0.06 0.4466$\pm$0.05 0.5095$\pm$0.08 0.4959$\pm$0.05 0.4959$\pm$0.05 0.5101$\pm$0.04 0.5426$\pm$0.02
TT 0.5569$\pm$0.05 0.5593$\pm$0.04 0.5563$\pm$0.04 0.5408$\pm$0.03 0.5386$\pm$0.03 0.5637$\pm$0.03 0.5640$\pm$0.04 0.5465$\pm$0.04 0.5471$\pm$0.04 0.5785$\pm$0.03
AC 0.7281$\pm$0.12 0.7783$\pm$0.08 0.6335$\pm$0.12 0.6354$\pm$0.12 0.7683$\pm$0.08 0.7488$\pm$0.14 0.7703$\pm$0.08 0.6762$\pm$0.14 0.6293$\pm$0.14 0.8206$\pm$0.08
VT 0.8625$\pm$0.01 0.8699$\pm$0.00 0.8699$\pm$0.00 0.8761$\pm$0.00 0.8667$\pm$0.00 0.8531$\pm$0.09 0.8736$\pm$0.00 0.8736$\pm$0.00 0.8713$\pm$0.00 0.8943$\pm$0.00
$\overline{\textit{AR}}$ 7.58 6.46 5.71 5.13 7.08 4.67 5.33 5.63 6.33 1.08
Table 4. Clustering performance evaluated by ARI (with range [-1,1]). “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.7638$\pm$0.18 0.7555$\pm$0.19 0.7232$\pm$0.19 0.7407$\pm$0.19 0.7307$\pm$0.22 0.7148$\pm$0.20 0.8650$\pm$0.22 0.7350$\pm$0.19 0.6968$\pm$0.17 0.9620$\pm$0.12
NS 0.0376$\pm$0.03 0.0341$\pm$0.02 0.0283$\pm$0.02 - 0.0393$\pm$0.03 0.0779$\pm$0.08 0.0309$\pm$0.03 0.0376$\pm$0.03 0.0376$\pm$0.03 0.1015$\pm$0.06
AP 0.0022$\pm$0.01 0.0171$\pm$0.02 0.0530$\pm$0.03 0.0322$\pm$0.02 0.0178$\pm$0.02 0.0355$\pm$0.00 0.0076$\pm$0.02 0.0150$\pm$0.02 0.0066$\pm$0.02 0.0524$\pm$0.00
DS 0.1077$\pm$0.13 0.1077$\pm$0.13 0.2347$\pm$0.20 0.2347$\pm$0.20 0.1844$\pm$0.20 0.1556$\pm$0.19 0.0512$\pm$0.05 0.1844$\pm$0.20 0.1921$\pm$0.19 0.2590$\pm$0.16
CS -0.0023$\pm$0.01 0.0088$\pm$0.04 0.0295$\pm$0.04 0.0250$\pm$0.03 0.0248$\pm$0.04 0.0647$\pm$0.03 0.0342$\pm$0.02 0.0234$\pm$0.03 0.0040$\pm$0.02 0.0627$\pm$0.03
HR -0.0081$\pm$0.01 -0.0008$\pm$0.01 -0.0013$\pm$0.01 -0.0010$\pm$0.01 -0.0046$\pm$0.01 -0.0033$\pm$0.01 -0.0149$\pm$0.00 0.0038$\pm$0.02 -0.0064$\pm$0.01 0.0368$\pm$0.01
ZO 0.5787$\pm$0.17 0.6085$\pm$0.17 0.5996$\pm$0.17 0.6091$\pm$0.15 0.6063$\pm$0.16 0.6094$\pm$0.12 0.6442$\pm$0.07 0.6138$\pm$0.15 0.6143$\pm$0.15 0.7536$\pm$0.11
BC 0.0046$\pm$0.02 0.0240$\pm$0.05 0.0811$\pm$0.07 0.0777$\pm$0.07 0.0128$\pm$0.04 0.0523$\pm$0.05 -0.0033$\pm$0.00 0.0605$\pm$0.08 0.0692$\pm$0.08 0.0799$\pm$0.03
LG 0.0843$\pm$0.05 0.0951$\pm$0.03 0.1027$\pm$0.06 0.1432$\pm$0.07 0.1018$\pm$0.04 0.1544$\pm$0.09 0.1715$\pm$0.04 0.1095$\pm$0.06 0.1262$\pm$0.05 0.1552$\pm$0.03
TT 0.0170$\pm$0.03 0.0184$\pm$0.02 0.0181$\pm$0.02 0.0095$\pm$0.01 0.0091$\pm$0.01 0.0156$\pm$0.01 0.0198$\pm$0.02 0.0120$\pm$0.02 0.0124$\pm$0.02 0.0226$\pm$0.02
AC 0.2596$\pm$0.18 0.3336$\pm$0.12 0.1210$\pm$0.15 0.1199$\pm$0.13 0.3092$\pm$0.11 0.3184$\pm$0.22 0.3107$\pm$0.11 0.1913$\pm$0.18 0.1349$\pm$0.18 0.4313$\pm$0.16
VT 0.5247$\pm$0.02 0.5463$\pm$0.01 0.5463$\pm$0.01 0.5648$\pm$0.01 0.5368$\pm$0.01 0.5200$\pm$0.18 0.5572$\pm$0.00 0.5572$\pm$0.00 0.5503$\pm$0.00 0.6207$\pm$0.00
$\overline{\textit{AR}}$ 7.88 5.92 5.50 5.38 6.71 5.08 5.29 5.50 6.42 1.33
4.2. Clustering Performance Evaluation
In this section, we investigate the clustering performance of different algorithms and statistically analyze the superiority of OCL.
4.2.1. Clustering performance on different datasets
The clustering results obtained by different algorithms on various datasets in terms of different validity indices are shown in Tables 3 and 4. The best and second-best results on each dataset are highlighted in bold and underlined, respectively. The observations include the following four aspects: (1) Overall, OCL reaches the best result on most datasets, indicating its superiority in clustering. (2) On the CA index, although OCL does not perform significantly better than the second-best method on the AP and TT datasets, the second-best methods are different on these datasets, but OCL has stable performance on them. (3) On the CA index, although OCL does not perform the best on CS dataset, its performance is close to the best result (the difference is less than 0.005), which still proves OCL’s competitiveness. (4) In terms of the ARI index, there are four datasets on which OCL does not perform the best, namely AP, CS, BC, and LG. However, similarly, the differences between OCL and the best-performing algorithms on AP, CS, and BC datasets are very small. In addition, the result of CBDM is not reported on the NS dataset because all the NS’s attributes are independent of each other, making the inter-attribute dependence-based CBDM fail in measuring distances.
4.2.2. Significance study
We first implement the Friedman (Friedman, 1937) test on the average ranks reported in the last rows in Tables 3 and 4. The corresponding p-values are 0.000017 and 0.000091 respectively, both passing the test under the confidence of 99% (p-value $=0.01$). On this basis, results of the two-tailed BD test (Demšar, 2006) with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 2, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively. According to the criterion presented in (Demšar, 2006), OCL is considered to perform significantly better than all the counterparts that fall outside the right boundary of the CD intervals. It can be seen that the proposed OCL is significantly superior to all counterparts.
Figure 2. Results of the two-tailed BD tests w.r.t. the CA (a) and ARI (b) performance in Tables 3 and  4, respectively.
4.3. Ablation Studies
To demonstrate the effectiveness of the core components of OCL and to verify our hypothesis on the orders of categorical data attributes, we design several variants of OCL for ablation studies from both algorithm and data perspectives.
From the algorithm aspect, to evaluate the effectiveness of the proposed Order Distance (OD), we compare OCL with its variant OCLI, which uses normalized equidistant order distance without considering the probability distribution of possible values. To evaluate the effectiveness of the Order Learning (OL) mechanism, we further remove the process of order learning from OCLI and only update the order once in the first iteration, which forms another variant OCLII. To completely discard Order Information (OI), we further make OCLII use only the traditional Hamming distance, thus forming the third variant OCLIII. Table 5 compares the clustering performance of OCL and the three variants.
It can be observed that the performance of OCL is superior to all variants. More specifically, OCL outperforms OCLI significantly on 11 out of all 12 datasets. This indicates that OD can effectively utilize the distribution information of the attribute values within different clusters for reasonable order distance measurement. Furthermore, the performance of OCLI is also significantly better than OCLII on all 12 datasets, which verifies the effectiveness of the OL mechanism in searching optimal orders during clustering. In addition, compared to the OCLIII variant that completely disregards OI, OCLII has better performance on 11 out of all 12 datasets, which proves that even a roughly estimated initial order can still significantly and extensively improve clustering performance.
Table 5. CA performance of ablated OCL variants formed by successively ablating the three core components, i.e., Order Distance (OD), Order Learning (OL), and Order Information (OI). The green arrows indicate the performance improvements of an OCL variant over its previous more basic variant. “$\overline{\textit{AR}}$” column reports the average performance rankings.
OCL Components Variants Datasets $\overline{\textit{AR}}$
OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9830$\uparrow$   0.3573$\uparrow$   0.6222$\uparrow$   0.7458$\uparrow$   0.6313   0.4326$\uparrow$   0.7792$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   0.5785$\uparrow$   0.8206$\uparrow$   0.8943$\uparrow$   1.17$\uparrow$
$\checkmark$ $\checkmark$ OCLI   0.9149$\uparrow$   0.3364$\uparrow$   0.6085$\uparrow$   0.7283$\uparrow$   0.6500$\uparrow$   0.3765$\uparrow$   0.7525$\uparrow$   0.5643$\uparrow$   0.4872$\uparrow$   0.5784$\uparrow$   0.7546$\uparrow$   0.8942$\uparrow$   2.00$\uparrow$
$\checkmark$ OCLII   0.8681$\uparrow$   0.3359$\uparrow$   0.6069$\uparrow$ 0.  7125$\uparrow$   0.6388$\uparrow$ 0.3712   0.7000$\uparrow$   0.5636$\uparrow$   0.4453$\uparrow$   0.5762$\uparrow$   0.7545$\uparrow$   0.8828$\uparrow$   2.92$\uparrow$
OCLIII 0.8468 0.3208 0.5392 0.6525 0.5463 0.3758 0.6624 0.5472 0.4277 0.5569 0.7281 0.8625 3.92
From the data aspect, to validate our hypothesis that correct order information is crucial to categorical data clustering, we conduct ablation experiments on datasets containing both nominal and ordinal attributes. We compare the proposed OCL, which learns orders for both the nominal and ordinal attributes, to two different variants. The first one is LNRO (Learn Nominal Reserve Ordinal), which retains the semantic order information of the original ordinal attribute and only learns orders for the nominal attributes. The other one is RNRO (Reserve Nominal Reserve Ordinal), which discards the order learning process for both nominal and ordinal attributes. Table 6 compares the clustering performance of the three variants. It can be observed that on five out of the six datasets, the performance of LNRO is superior to RNRO. This verifies our hypothesis that the nominal attributes also contain potential order information that is effective for clustering. Significantly, we can find that the OCL method outperforms LNRO on five out of the six datasets. This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering than the original semantic orders.
Table 6. CA performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
Learn Order for Variants Datasets $\overline{\textit{AR}}$
OA NA NS AP CS HR BC LG
OCL   0.3573$\uparrow$   0.6222$\uparrow$ 0.6313   0.4326$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   1.17$\uparrow$
LNRO 0.3281   0.6074$\uparrow$   0.6350$\uparrow$   0.3985$\uparrow$   0.5476$\uparrow$   0.4818$\uparrow$   2.00$\uparrow$
RNRO 0.3308 0.5984 0.6225 0.3902 0.5276 0.4493 2.83
4.4. Convergence and Efficiency Evaluation
To evaluate the convergence of OCL, we execute it on all 12 datasets and present the convergence curves in Figure 3. The horizontal and vertical axes represent the number of learning iterations and the value of the objective function $L$, respectively. In each sub-figure of Figure 3, the blue triangles mark the iterations of order updates, while the red square indicates the iteration that OCL converges. It can be observed that, after each order update, the value of $L$ further decreases. This indicates that the proposed order learning mechanism is consistent with the optimization of the clustering objective function. Moreover, for all the tested datasets, OCL completes the learning process within 30 iterations with at most three updates of the orders. This reflects that OCL converges efficiently with guarantee, which conforms to the analysis in Section 3.5.
Figure 3. Convergence curve of OCL on all 12 datasets.
Furthermore, to evaluate the execution efficiency of OCL, we construct large synthetic datasets with different numbers of samples and attributes to investigate the execution time of OCL in comparison with other counterparts. The results reported in Figure 4 show that the execution time of OCL is better than UDMC and DLC on large-scale datasets, and better than H2H, UDMC, and CBDM on high-dimensional datasets. Additionally, as the number of samples or attributes increases, the execution time of OCL increases linearly, which is consistent with the time complexity analysis in Section 3.5. In summary, OCL is efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
Figure 4. Execution time on synthetic datasets with different numbers of samples $n$ and attributes $d$.
4.5. Visualization of Clustering Effect
To demonstrate the cluster discrimination ability of the OCL method, we utilize the distance between attribute values learned by OCL, CBDM, and ADC respectively to encode the attributes of the VT dataset. Subsequently, the t-SNE (Maaten and Hinton, 2008) is introduced to process the encoded data into two-dimensional space and visualize it by marking the samples with real labels in different colors. It can be seen from Figure 5 that OCL has significantly better cluster discrimination ability than CBDM and ADC.
(a) CBDM @@@@ VT
(b) ADC @@@@ VT
(c) OCL @@@@ VT
Figure 5. t-SNE visualization of the VT dataset represented by CBDM, ADC, and OCL.
5. Concluding Remarks
This paper introduces a new finding that knowing the correct order relation of attribute values is critical to cluster analysis of categorical data. Accordingly, we propose the OCL paradigm that allows joint learning of data partitions and order relation of attribute values for accurate clustering. The proposed learning mechanism can infer a more appropriate order among categorical attribute values based on the current clusters, thus yielding more compact clusters based on the updated order. The data partition and order learning are iteratively executed to approach the clustering objective, i.e., maximizing the overall similarity of intra-cluster samples. As a result, more accurate clustering results and insightful orders can be obtained simultaneously. OCL converges quickly with guarantee, and extensive experiments conducted on 12 real benchmark datasets fully demonstrate its superiority in comparison with state-of-the-art clustering methods. A case study has also been provided to show the rationality of the learned order in understanding unlabeled categorical data.
The proposed OCL method is also not exempt from limitations. Although it demonstrates the potential to enhance the performance of mixed data clustering, the deep connection between the treatment of categorical and numerical attributes remains to be explored. In addition, it would also be promising to extend OCL to more complex datasets with non-static distribution, an unknown number of clusters, and extremely imbalanced clusters.

### v3

@@
 Order Is All You Need for Categorical Data Clustering
-Abstract.
-Categorical data composed of nominal valued attributes are ubiquitous in knowledge discovery and data mining tasks.
-Due to the lack of well-defined metric space, categorical data distributions are difficult to intuitively understand.
-Clustering is a popular technique suitable for data analysis.
-However, the success of clustering often relies on reasonable distance metrics, which happens to be what categorical data naturally lack.
-Therefore, the cluster analysis of categorical data is considered a critical but challenging problem.
-This paper introduces the new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding the categorical data clusters.
-To automatically obtain the orders, we propose a new learning paradigm that allows joint learning of clusters and the orders.
-It turns out that clustering with order learning achieves superior clustering accuracy, and the learned orders provide intuition for understanding the cluster distribution of categorical data.
-Extensive experiments with statistical evidence and case studies have verified the effectiveness of the new “order is all you need” insight and the proposed method.
-The source code of this submission is temporarily opened [^1] https://anonymous.4open.science/r/OCL-demo.
+Abstract
+Categorical data composed of qualitative valued attributes are ubiquitous in machine learning tasks.
+Due to the lack of a well-defined metric space, categorical data distributions are difficult to understand intuitively.
+Clustering is a popular data analysis technique suitable for understanding data distribution.
+However, the success of clustering often relies on reasonable distance metrics, which happen to be what categorical data naturally lack.
+This paper, therefore, introduces a new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding categorical data clusters, because the essence of clustering is to order the clusters in terms of their admission to samples.
+To obtain the clustering-friendly order distance metric, we propose a new learning paradigm that allows joint learning of clusters and the orders.
+It alternatively partitions data into clusters based on order distance, and estimates the most likely orders according to the clusters.
+The algorithm achieves superior clustering accuracy with a convergence guarantee, and the learned orders facilitate the understanding of the non-intuitive categorical data from the cluster perspective.
+Experiments with ablation studies, significance tests, case studies, etc., have validated the efficacy of the proposed method.
+The source code is available at https://anonymous.4open.science/r/OCL-demo.
 Cluster analysis, Categorical data, Order relationship, Distance learning, Partitional clustering
-submissionid: 51ccs: Computing methodologies Cluster analysisccs: Computing methodologies Learning latent representationsccs: Computing methodologies Mixture modeling
-1. Introduction
+1 Introduction
+Figure 1: Clustering accuracy of $k$-modes by treating 12 real benchmark datasets with Randomly generated Orders (RO) (color-filled regions), without orders (green bars), and with inherent semantic orders (brown bars).
+It can be seen that the performance upper bound of color-filled regions is obviously higher than that of the bars, indicating that even random orders can easily achieve significant accuracy improvements.
+Detailed statistics of the datasets and settings are provided in Table 6 and Section A.3.1 of the Appendix.
 […]
-In the literature, there are two main types of clustering algorithms for categorical data: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia and Cheung, 2018) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004) that adjust the sample-cluster affiliations based on certain cluster quality measures.
+In the literature11footnotetext: Data composed of nominal and ordinal attributes., clustering algorithms for categorical data can be roughly classified into: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia & Cheung, 2018)(Bai & Liang, 2022) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004)(Cai et al., 2024)(Zhang et al., 2025) that adjust the sample-cluster affiliations based on certain cluster quality measures.
 Their performance is commonly dominated by the metric space represented by the adopted measures.
-Thus, the focus of subsequent research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Naouali et al., 2020).
-[…]
+Thus, the focus of current research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Alamuri et al., 2014).
 The conventional Hamming distance measure and one-hot encoding strategy are very common in categorical data clustering, but they both simply represent the distance between two attribute values based on whether the two values are identical or not, which is far from informative.
-Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler and Lindenbaum, 2011), inter-attribute relationship (Ahmad and Dey, 2007)(Le and Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang and Cheung, 2022b).
-On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed Although all the above advanced techniques have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting different clustering tasks on various categorical datasets.
+Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler & Lindenbaum, 2011), inter-attribute relationship (Ahmad & Dey, 2007)(Le & Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang & Cheung, 2022).
+On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015) (Zhang et al., 2023), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed.
+Although they have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting to different clustering tasks, i.e., $T(k,X)$ subject to different numbers of clusters $k$ on various categorical datasets $X$.
 […]
@@
 The overall order is formed based on the similarity orders of the clusters to the sample reflected by different attributes.
-Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of values, which makes it easy to obtain the value order.
+Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of its values, which makes it easy to obtain the value order.
 According to the above analysis, we also note that although most existing categorical data clustering approaches do not specify the order of attribute values, their clustering process still passively forms the order through similarity measurement, regardless of whether there is an obvious semantic order of the attribute values.
-Therefore, knowing the correct order of attribute values may be beneficial to categorical data clustering, and the experimental results in Figure 1 confirmed this.
+Therefore, knowing the correct order of attribute values may considerably benefit categorical data clustering, and the demo experimental results in Figure 1 confirmed this.
 It can be seen that even randomly generated orders can easily achieve significant accuracy improvements.
 Another interesting observation is that, although attribute values of some datasets (i.e., the right eight datasets) are with obvious semantic orders, they are rarely optimal for clustering.
-This is because the semantic orders of attributes are relatively independent of specific clustering tasks, and this information gap can prevent some semantic orders from well-serving clustering.
+This is because the human-assigned semantic orders are relatively subjective and may thus be independent of specific clustering tasks $T(k,X)$, preventing the semantic orders from well-serving clustering.
 Therefore, obtaining the optimal orders w.r.t.
-certain clustering tasks is the key to achieving accurate categorical data clustering.
+certain clustering tasks is the key to accurate clustering.
 This paper, therefore, proposes a new clustering paradigm that dynamically learns orders based on the current data partition during clustering.
-Since the order learning is connected to the clustering task, the optimal order for clustering can thus be automatically obtained.
+Since we connect the order learning to the clustering task, the optimal order for clustering can be obtained.
 To make full use of the learned orders, a distance learning mechanism is also introduced to dynamically measure the sample-cluster affiliation under the guidance of the orders.
 It turns out that the learned orders can significantly improve clustering accuracy and can also intuitively help understand the distribution and clustering effect of categorical data.
-Extensive experimental evaluations demonstrate the superiority of the proposed method.
-The main contributions can be summarized into three points:
-  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed for the first time.
-Accordingly, a new clustering paradigm with order distance learning is proposed.
-  • We propose a distance structure learning mechanism to automatically obtain the order and distance.
-This learning process is jointly performed with clustering, making the clustering performance more robust to various datasets.
-  • Both the learning process and the learned order are highly interpretable, which is helpful for understanding complex categorical data.
-Moreover, the algorithm is efficient, strictly convergent, and can be easily extended to mixed data [^2] Dataset composed of a mixture of categorical and numerical attributes..
-2. Related Work
-This section overviews the related works in data encoding, representation learning, and consideration of value order, for categorical data clustering.
+Extensive experiments demonstrate the superiority of the proposed paradigm.
+The main contributions of this work are summarized into three points:
+  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed.
+Accordingly, a new clustering paradigm with order distance learning is presented.
+  • We propose an iterative learning mechanism to realize the optimization of the order and distance.
+This learning process is jointly performed with clustering, making the clustering performance more adaptive to various clustering scenarios $T(k,X)$.
+  • Both the learning process and learned orders are highly interpretable, which helps understand complex categorical data.
+Moreover, the algorithm is efficient, strictly convergent, and can be easily adapted for mixed data [^2] Dataset composed of categorical and numerical attributes.
+Corresponding evaluation can be found in Appendix A.10..
+2 Related Work
+This section overviews the related works in: 1) data encoding, 2) representation learning, and 3) consideration of value order, for categorical data clustering.
 The data encoding techniques convert the categorical data into numerical values based on certain strategies, and then perform clustering on the numerical encoding.
-The conventional one-hot encoding and Hamming distance construct similar distance structures for categorical data by assigning distances “1” and “0” to any pair of different and identical values, respectively.
-As they cannot distinguish different dissimilarity degrees (dos Santos and Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable during clustering.
-Some statistic-based distance measures (Lin, 1998)(Cheung and Jia, 2013)(dos Santos and Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples.
-They take a common principle that two similar values may have more consistent statistics within the corresponding attribute or cluster.
-Some metrics (Ahmad and Dey, 2007)(Le and Ho, 2005) further define distance considering the relevance between attributes.
-The representative measure proposed in (Ienco et al., 2012) selects more relevant attributes for the distance measurement.
-Later, the measures (Jia et al., 2016)(Jian et al., 2018a) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have been presented.
-Coupling-based encoding strategies (Jian et al., 2018c)(Jian et al., 2017) have also been designed accordingly to encode the couplings.
-All the above techniques perform value-level encoding to reflect sample-level distances, while the work (Qian et al., 2015) directly encodes a data sample by concatenating the distances from it to all the other samples.
+As the conventional one-hot encoding and Hamming distance cannot distinguish different dissimilarity degrees (dos Santos & Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable.
+Some statistic-based distance measures (Lin, 1998)(Cheung & Jia, 2013)(dos Santos & Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples.
+Some metrics (Ahmad & Dey, 2007)(Le & Ho, 2005)(Ienco et al., 2012) further define distance considering the relevance between attributes.
+Later, the measures (Jia et al., 2016)(Jian et al., 2018a) and encoding strategies (Jian et al., 2018c)(Jian et al., 2017)(Qian et al., 2015) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have also been presented.
 Recently, more advanced representation learning methods have been proposed to learn the categorical data encoding.
 The work proposed in (Zhu et al., 2022) projects dataset through different kernels for more informative coupling representation, and then learns the kernel parameters to obtain task-oriented representation.
+In contrast, (Cai et al., 2024) reveals the complex nested multi-granular cluster effect to obtain multi-granular distribution encoding representation of categorical data.
 Some approaches (Jian et al., 2018b)(Zhu et al., 2020) have also been designed to simultaneously encode the numerical and categorical attributes.
 However, the involved hyper-parameters have a great impact on their performance, and the appropriate parameter selection will be laborious in different real applications.
-Although a parameter-free learning approach (Jia and Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
+Although a parameter-free learning approach (Jia & Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
 Some new advances proposed with the consideration of value order further distinguish categorical attributes into nominal and ordinal attributes.
-The distance metrics (Zhang et al., 2020)(Zhang and Cheung, 2022b) define the distance between attribute values based on the data statistical information with considering the semantic order.
-Later, a representation learning approach was introduced by (Zhang and Cheung, 2020), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts.
-Later, the works (Zhang and Cheung, 2022a)(Zhang and Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering.
-The most recent work (Zhang et al., 2022) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering.
+The distance metrics (Zhang et al., 2020b)(Zhang & Cheung, 2022) define the distance between attribute values based on the data statistical information with considering the semantic order.
+Later, a representation learning approach was introduced by (Zhang et al., 2020a), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts.
+Later, the works (Zhang et al., 2022b)(Zhang & Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering.
+The most recent work (Zhang et al., 2022a) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering.
 The above recent advances achieve considerable clustering performance improvements and empirically prove the importance of value order in categorical data clustering.
@@
 Moreover, they directly adopt this order for clustering, and overlook that the semantic order may not be suitable under different clustering scenarios as shown in Figure 1.
-3. Proposed Method
-This section first formulates the problem, then proposes the order distance definition and order learning mechanism, and finally introduces the joint learning of the orders and clusters.
-Analyses of time complexity and convergence are also provided.
-3.1. Problem Formulation
+3 Proposed Method
+This section first formulates the problem, then proposes the order distance definition, order learning mechanism, and the joint learning of the order-guided distances and clusters with time complexity and convergence analysis.
+3.1 Problem Formulation
 The problem of categorical data clustering with order learning is formulated as follows.
-Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples (e.g., a collection of $n$ bank clients).
-Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$ (e.g., occupation, gender, priority, etc.).
-Each attribute $\mathbf{a}_{r}$ (e.g., occupation) can be denoted as a column vector $\mathbf{a}_{r}=[x_{1,r},x_{2,r},...,x_{n,r}]$ composed of the occupation description of all the $n$ clients.
-For the categorical attribute “occupation” $\mathbf{a}_{r}$, all its $n$ values are taken from a limited number of qualitative possible values $V_{r}=\{v_{r,1},v_{r,2},...,v_{r,l_{r}}\}$, e.g., $\{\text{lawyer},\text{doctor},...,\text{scientist}\}$.
-The subscripts indicate the sequential number of an attribute and the sequential number of the attribute’s possible value.
-For example, doctor is the second value in the $r$-th attribute “occupation”, so we have $v_{r,2}=\text{doctor}$.
+Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples.
+Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$.
 We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data.
-Please note that the sequential numbers of possible values only distinguish them, and do not indicate their order.
 Assume that there is an optimal order indicating an appropriate underlying distance structure of possible values w.r.t.
-a clustering task.
+each clustering task $T(k,X)$ subject to the sough number of clusters $k$ specified by the users and the characteristics of the dataset $X$.
 We denote the optimal order of an attribute $\mathbf{a}_{r}$ as $O^{*}_{r}=\{o^{*}(v_{r,1}),o^{*}(v_{r,2}),...,o^{*}(v_{r,l_{r}})\}$ where the superscript “*” marks the optimum and the value of $o^{*}(v_{r,g})$ is an integer reflecting the unique ranking of $v_{r,g}$ among all the $l_{r}$ possible values in $V_{r}$.
-By still taking the “occupation” attribute as an example, if its three possible values $v_{r,1}=\text{lawyer}$, $v_{r,2}=\text{doctor}$, and $v_{r,3}=\text{scientist}$, ranking second (i.e., $o(\text{lawyer})=2$), third (i.e., $o(\text{doctor})=3$), and first (i.e., $o(\text{scientist})=1$), respectively.
-Then the corresponding order of these values are $O_{r}=\{o(\text{lawyer}),o(\text{doctor}),o(\text{scientist})\}=\{2,3,1\}$.
-Such order reflects a distance structure that roughly satisfies $\text{dist}(v_{r,g},v_{r,h})\propto|o(v_{r,g})-o(v_{r,h})|$.
-More specifically, the following distance relationship should hold:
-  1. (1)
-    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{doctor},\text{lawyer})$,
-  2. (2)
-    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{lawyer},\text{scientist})$,
-because “doctor” and “scientist” are further away in order, compared to “lawyer” and “doctor” that are adjacent in order.
 We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$.
 Then the order information reflecting the distance structure can guide the partition of data objects during clustering.
-Our approach, i.e., learning order relations, is more generative than most existing approaches that directly define the distance between attribute values.
-We focus more on how the intrinsic relationship between values affects the distance between samples in forming clusters, rather than simply defining the clustering-irrelevant distance between values based on data statistics, e.g., occurrence probability of values in an attribute.
-For a crisp partitional clustering task to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
-(1) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$
+For a crisp partitional clustering task $T(k,X)$ to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
+$L(\mathbf{Q},O)=\mathop{\sum\limits_{m=1}^{k}{\mathop{\sum}\limits_{i=1}^{n}}}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$ (1)
 where $\mathbf{Q}$ is an $n\times k$ matrix with its $(i,m)$-th entry $q_{i,m}$ indicating the affiliation between sample $\mathbf{x}_{i}$ and cluster $C_{m}$.
 Specifically, $q_{i,m}=1$ indicates that $\mathbf{x}_{i}$ belongs to $C_{m}$, while $q_{i,m}=0$ indicates that $\mathbf{x}_{i}$ belongs to a cluster other than $C_{m}$, which can be written as:
-(2) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$
+$q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$ (2)
 with $\sum_{i=m}^{k}q_{i,m}=1$.
-The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$.
-Therefore, the problem to be solved can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$.
-When the algorithm converges, a sub-optimal solution with $k$ compact bank client clusters is expected to be obtained.
-In general, “compact” means that the bank clients with more similar characteristics (i.e., with more similar value descriptions across all the attributes) are gathered into the same cluster, and different clusters are relatively distinguishable.
-Frequently used symbols in this paper and the corresponding explanations are sorted out in Table 1 below.
-Then we discuss how to define the order distance and learn order for accurate categorical data clustering.
-Table 1. Frequently used symbols. Note that we uniformly use lowercase, uppercase, bold lowercase, and bold uppercase to indicate value, set, vector, and matrix, respectively.
-Symbol Explanation
-$X$, $A$, $O$, and $C$ Dataset, attribute set, order set, and cluster set
-$\mathbf{x}_{i}$ and $x_{i,r}$ $i$-th data object and $r$-th value of $\mathbf{x}_{i}$
-$\mathbf{a}_{r}$, $V_{r}$, and $v_{r,g}$ $r$-th attribute, its value set, and $g$-th value of $V_{r}$
-$l_{r}$ Total number of possible values in $V_{r}$
-$O_{r}$ Order values (integers) corresponding to $V_{r}$
-$o(v_{r,g})$ Order value (integer) of possible value $v_{r,g}$
-$\mathbf{Q}$ and $C_{m}$ Object-cluster affiliation matrix and $m$-th cluster
-$q_{i,m}$ $(i,m)$-th entry of $\mathbf{Q}$ indicating affiliation of $\mathbf{x}_{i}$ to $C_{m}$
-$\Theta(\mathbf{x}_{i},C_{m};O)$ Overall order distance between $\mathbf{x}_{i}$ and $C_{m}$
-$\theta(x_{i,r},\textbf{p}_{m,r})$ Order distance between $\mathbf{x}_{i}$ and $C_{m}$ reflected by $\mathbf{a}_{r}$
-$\mathbf{D}$ and $\mathbf{d}_{i,r}$ Order distance matrix and its $(i,r)$-th entry
-$d_{i,r,g}$ Order distance between $x_{i,r}$ and $v_{r,g}$ ($g$-th value of $\mathbf{d}_{i,r}$)
-$\mathbf{P}$ and $\mathbf{p}_{m,r}$ Conditional probability matrix and its $(m,r)$-th entry
-$p_{m,r,g}$ Occurrence probability of $v_{r,g}$ in $C_{m}$ ($g$-th value of $\mathbf{p}_{m,r}$)
-$O_{r,m}$ Order values of $O_{r}$ obtained through $C_{m}$
-$L_{r,m}$ A fraction of $L$ jointly contributed by $\mathbf{a}_{r}$ and $C_{m}$
-3.2. Order Distance
-We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective.
-The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ is defined as:
-(3) $\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$
+$\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity (also called distance hereinafter) between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$.
+Therefore, the problem can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$.
+Frequently used symbols and more detailed problem settings with intuitive examples are provided in Appendix A.1.
+Then we present the definition of order distance $\Theta$ subject to order $O$ and discuss how to learn $O$ for categorical data clustering [^3] In the following, $\Theta$ is expressed in the form of a matrix $\mathbf{D}$ to facilitate the description of the distance definition and computation..
+3.2 Order Distance
+We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective:
+$\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$ (3)
 where
-(4) $\theta(x_{i,r},\textbf{p}_{m,r})=\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}.$
+$\theta(x_{i,r},\textbf{p}_{m,r})=\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}$ (4)
 is the order distance between $\mathbf{x}_{i}$ and all the samples in $C_{m}$ reflected by the $r$-th attribute.
 $\mathbf{d}_{i,r}$ is an $l_{r}$-dimensional vector containing the order distances between $x_{i,r}$ and possible values in $V_{r}$, and $\mathbf{p}_{m,r}$ is an $l_{r}$-dimensional vector describing the probability distribution of $V_{r}$ within cluster $C_{m}$, which can be written as:
-(5) $\mathbf{d}_{i,r}=[d_{i,r,1},d_{i,r,2},...,d_{i,r,l_{r}}]$
+$\mathbf{d}_{i,r}=[d_{i,r,1},d_{i,r,2},...,d_{i,r,l_{r}}]$ (5)
 and
-(6) $\mathbf{p}_{m,r}=[p_{m,r,1},p_{m,r,2},...,p_{m,r,l_{r}}],$
+$\mathbf{p}_{m,r}=[p_{m,r,1},p_{m,r,2},...,p_{m,r,l_{r}}],$ (6)
 respectively. The $g$-th value of $\mathbf{d}_{i,r}$ is defined as the normalized order distance between $x_{i,r}$ and $v_{r,g}$, which can be written as:
-(7) $d_{i,r,g}=\frac{|o(x_{i,r})-o(v_{r,g})|}{l_{r}-1}$
+$d_{i,r,g}=\frac{|o(x_{i,r})-o(v_{r,g})|}{l_{r}-1}$ (7)
 where $l_{r}-1$ in the denominator is to normalize the order distance into the interval $[0,1]$ to ensure distance comparability of different attributes.
 The $g$-th value of $\mathbf{p}_{m,r}$ is the occurrence probability of $v_{r,g}$ in cluster $C_{m}$, which can be written as:
-(8) $p_{m,r,g}=\frac{card(X_{r,g}\cap C_{m})}{card(C_{m})}$
+$p_{m,r,g}=\frac{card(X_{r,g}\cap C_{m})}{card(C_{m})}$ (8)
 where $X_{r,g}=\{\mathbf{x}_{i}|x_{i,r}=v_{r,g}\}$ is a sample set collecting all the samples in $X$ with their $r$-th values equal to $v_{r,g}$, and the cardinality function $card(\cdot)$ counts the number of elements in a set.
-Remark 1.
+Remark 3.1.
 Rationality of the order distance.
@@
 By making the distance $\mathbf{d}_{i,r}$ dependent on the order $O$ through Eq.
-(7), we facilitate a direct connection between the order learning and the clustering tasks.
-Such a connection leverages the optimization of the order $O$ for more accurate clustering in Section 3.4.
-3.3. Order Learning
+(7), a direct connection between the order learning and the clustering tasks is facilitated leveraging the optimization of the order $O$ for more accurate clustering in Section 3.4.
+3.3 Order Learning
 Based on the above-defined order distance, we discuss how to obtain the optimal order:
-(9) $O^{*}=\arg\min_{O}L(\mathbf{Q},O).$
-A simple but extremely laborious solution is to search different order combinations across all the attributes.
+$O^{*}=\arg\min_{O}L(\mathbf{Q},O).$ (9)
+A simple but laborious solution is to search different order combinations across all the attributes.
 There are a total of $\prod_{r=1}^{d}l_{r}!$ combinations, and each combination should be evaluated on $n$ samples by accumulating their sample-cluster distances calculated by Eq.
@@
 Accordingly, optimal order of $\mathbf{a}_{r}$ w.r.t.
-$C_{m}$ denoted as $O^{*}_{r,m}=\{o^{*}_{m}(v_{r,1}),o^{*}_{m}(v_{r,2}),...,o^{*}_{m}(v_{r,l_{r}})\}$ can be obtained through the efficient intra-cluster traversal search:
-(10) $O^{*}_{r,m}=\arg\min_{O_{r}}L_{r,m}$
+$C_{m}$ denoted as $O^{*}_{r,m}=\{o^{*}_{m}(v_{r,1}),o^{*}_{m}(v_{r,2}),...,o^{*}_{m}(v_{r,l_{r}})\}$ can be obtained through efficient intra-cluster traversal search:
+$O^{*}_{r,m}=\arg\min_{O_{r}}L_{r,m}$ (10)
 where $L_{r,m}$ is an inner sum component of the objective $L$.
@@
 (1) and (3) as:
-(11) $L_{r,m}=\sum_{\mathbf{x}_{i}\in C_{m}}\theta(x_{i,r},\textbf{p}_{m,r}).$
+$L_{r,m}=\sum_{\mathbf{x}_{i}\in C_{m}}\theta(x_{i,r},\textbf{p}_{m,r}).$ (11)
 By searching for the optimal ranking in each cluster and then performing a linear combination based on the cluster size, the overall ranking of each possible value $v_{r,g}$ can be obtained by:
-(12) $\bar{o}(v_{r,g})=\sum_{m=1}^{k}\left(o^{*}_{m}(v_{r,g})\cdot\frac{card(C_{m})}{n}\right).$
+$\bar{o}(v_{r,g})=\sum_{m=1}^{k}\left(o^{*}_{m}(v_{r,g})\cdot\frac{card(C_{m})}{n}\right).$ (12)
 Since the value of $\bar{o}(v_{r,g})$ is not necessarily an integer, we sort the obtained rankings $\{\bar{o}(v_{r,1}),\bar{o}(v_{r,2}),...,\bar{o}(v_{r,l_{r}})\}$ of an attribute in ascending order to obtain the final ranking of values in $V_{r}$, and the corresponding integer order $\tilde{O}_{r}=\{\tilde{o}(v_{r,1}),\tilde{o}(v_{r,2}),...,\tilde{o}(v_{r,l_{r}})\}$ is obtained.
@@
 (10) - (12) as $\tilde{O}=\{\tilde{O}_{1},\tilde{O}_{2},...,\tilde{O}_{d}\}$.
-Remark 2.
+Remark 3.2.
 Principle of the order learning.
@@
 Through Eq.
-(11), the specific distribution of samples within each cluster is fully considered (see Remark 1) to appropriately reflect the component $L_{r,m}$ of the objective $L$.
+(11), the specific distribution of samples within each cluster is fully considered (see Remark 3.1) to appropriately reflect the component $L_{r,m}$ of the objective $L$.
 Hence, an expected minimization on a series of $L_{r,m}$s can be achieved by the determined orders.
@@
 Here, the cluster sizes are taken into account to accommodate more to larger clusters, thereby achieving a more effective reduction on $L$.
-3.4. Joint Order and Cluster Learning
+3.4 Joint Order Distance and Cluster Learning
 We facilitate joint learning of clusters and orders by integrating the order learning into the clustering objective, which can be rewritten in a more detailed form based on Eqs. (1), (3), and (4):
-(13) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}$
+$L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}$ (13)
 where the order distance $\mathbf{D}=\{\mathbf{d}_{i,r}|i\in\{1,2,...,n\},r\in\{1,2,...,d\}\}$ is dependent to the order $O$ according to Eqs.
@@
 Specifically, given $\hat{O}$, the order distance $\mathbf{D}$ can be updated accordingly, and then a new $\mathbf{Q}$ can be computed by:
-(14) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{y,r}\\ 0,&\text{otherwise}\\ \end{array}\right.$
+$q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{y,r}\\ 0,&\text{otherwise}\\ \end{array}\right.$ (14)
 with $i=\{1,2,...,n\}$ and $m=\{1,2,...,k\}$.
@@
 (10) - (12), and update $\mathbf{D}$ accordingly.
-In summary, $L$ is minimized by iteratively solving the following two sub-problems:
-  1. (1)
-    Fix $\hat{\mathbf{Q}}$ and $\hat{\mathbf{P}}$, solve the minimization problem $L(\hat{\mathbf{Q}},O)$ by searching $O$ and updating $\mathbf{D}$ accordingly;
-  2. (2)
-    Fix $\hat{O}$ and $\hat{\mathbf{D}}$, solve the minimization problem $L(\mathbf{Q},\hat{O})$ by iteratively computing $\mathbf{Q}$, and updating $\mathbf{P}$ according to $\mathbf{Q}$, until convergence;
-The whole algorithm named Order and Cluster Learning (OCL) is summarized as Algorithm 1. The next subsection provides analyses of time complexity and convergence to further explain OCL.
+In summary, $L$ is minimized by iteratively solving the following two sub-problems: 1) fix $\hat{\mathbf{Q}}$ and $\hat{\mathbf{P}}$, solve the minimization problem $L(\hat{\mathbf{Q}},O)$ by searching $O$ and updating $\mathbf{D}$ accordingly, and 2) fix $\hat{O}$ and $\hat{\mathbf{D}}$, solve the minimization problem $L(\mathbf{Q},\hat{O})$ by iteratively computing $\mathbf{Q}$, and updating $\mathbf{P}$ according to $\mathbf{Q}$, until convergence.
+The whole algorithm named Order and Cluster Learning (OCL) is summarized as Algorithm 1.
 Algorithm 1 OCL: Order and Cluster Learning Algorithm
@@
 13:  end while
-3.5. Complexity and Convergence Analysis
+3.5 Complexity and Convergence Analysis
 We analyze the time complexity to show the scalability of OCL, and then provide its convergence analysis.
-Theorem 1.
+Due to space limitation, we briefly summarize the analyses and refer the readers to Appendix A.2 for detailed proofs.
+Theorem 3.3.
+The time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$.
+Proof. The time complexity of $I$ inner iterations is $O(Inkd\varsigma)$, and the complexity of $E$ outer iterations, which is the overall complexity, is $O(EInkd\varsigma+Edkn\varsigma!)$.
+Theorem 3.4.
+OCL algorithm converges to a local minimum in a finite number of iterations.
+Proof. We prove the convergence of the inner loop (i.e., lines 4 - 9) and outer loop (i.e., lines 2 - 13) of Algorithm 1, respectively, with the following two Lemmas.
+Lemma 3.5.
+The inner loop converges to a local minimum of $L(\mathbf{Q},\hat{O})$ in a finite number of iterations with fixed $\hat{O}$.
+Lemma 3.6.
+The outer loop converges to a local minimum of $L(\mathbf{Q},O)$ in a finite number of iterations.
+Since both the inner and outer loop of Algorithm 1 converge to a local minimum, the result follows. □
+4 Experiments
+4.1 Experimental Settings
+Four types of experiments have been conducted to evaluate the proposed method.
+A brief outline of the experiments is given below: (1) Clustering performance evaluation with significance test (Section 4.2): We compared the proposed OCL with conventional and state-of-the-art methods on different types of datasets.
+To statistically analyze the performance differences between variant approaches, we also conducted a significance test on the clustering results; (2) Ablation studies (Section 4.3): To more specifically demonstrate the effectiveness of the core components of OCL and investigate the role of learned order in improving clustering performance, we conduct ablation studies from the algorithm aspect and data aspect.
+(3) Convergence and efficiency evaluation (Section 4.4): The values of the objective function after each learning iteration have been recorded to study the convergence of OCL.
+Its efficiency has also been studied with experiments on large datasets with scaling sizes of samples $n$ and attributes $d$.
+(4) Visualization of clustering effect (Section 4.5): To visually demonstrate the effectiveness of the order distance metric learned by OCL, we use it and the counterparts to numerically encode the data, and utilize the t-SNE for visualization.
+Eleven counterparts have been selected for the comparative study.
+The counterparts include traditional K-MoDes (KMD) algorithm (Huang, 1998), popular entropy-based metric LSM (Lin, 1998), and context-based distance measurement CBDM (Ienco et al., 2009).
+The latter two are representative measures proposed for categorical data, both of which can be embedded into KMD to conduct clustering.
+Besides, some state-of-the-art methods have also been taken into account, including Jia’s Distance Metric (JDM) (Jia et al., 2016), Unified Distance Metric based Clustering (UDMC) (Zhang & Cheung, 2022), HD based Clustering (HDC) (Zhang et al., 2022b), projection based clustering HARR (Zhang et al., 2022a), Distance Learning based Clustering (DLC) algorithm (Zhang et al., 2020a), graph metric based clustering algorithm ADC (Zhang & Cheung, 2023), Multi-granular-guided Categorical Data Clustering (MCDC) (Cai et al., 2024), an Adaptive Micro Partition and Hierarchical Merging clustering algorithm (AMPHM) (Zhang et al., 2025).
+Please note that we have set the parameters (if any) of these methods to the values recommended in the corresponding paper.
+In the experiments, each clustering result was recorded based on the average performance of 10 runs.
+Three validity indices, i.e., clustering accuracy (CA) (He et al., 2005), ARI coefficients (Rand, 1971)(Gates & Ahn, 2017), and standardized mutual information NMI (Estévez et al., 2009), have been selected to evaluate the clustering performance.
+Usually, larger values of these indices indicate better clustering performance.
+We also conduct the Bonferroni Dunn (BD) test (Demšar, 2006) on the performance of different methods and calculate the critical difference (CD) interval to statistically analyze the superiority of our OCL.
+Table 1: Clustering performance evaluated by CA. “$\overline{\textit{AR}}$” row reports the average performance rankings.
+Datasets KMD LSM JDM CBDM UDMC DLC HARR HDC ADC MCDC AMPHM OCL
+SB 0.8468$\pm$0.15 0.8106$\pm$0.16 0.7894$\pm$0.14 0.7957$\pm$0.15 0.7915$\pm$0.18 0.8000$\pm$0.14 0.9234$\pm$0.12 0.7915$\pm$0.15 0.7660$\pm$0.13 0.5851$\pm$0.12 0.7021$\pm$ 0.00 0.9830$\pm$0.05
+NS 0.3208$\pm$0.05 0.2975$\pm$0.04 0.2980$\pm$0.04 -- 0.3074$\pm$0.05 0.3386$\pm$0.07 0.3113$\pm$0.05 0.3208$\pm$0.05 0.3208$\pm$0.05 0.3370$\pm$0.01 -- 0.3573$\pm$0.05
+AP 0.5392$\pm$0.03 0.5656$\pm$0.04 0.6159$\pm$0.04 0.5926$\pm$0.03 0.5698$\pm$0.03 0.6032$\pm$0.00 0.5603$\pm$0.03 0.5608$\pm$0.04 0.5529$\pm$0.04 0.5566$\pm$0.03 0.5026$\pm$ 0.00 0.6222$\pm$0.00
+DS 0.6525$\pm$0.08 0.6525$\pm$0.08 0.7208$\pm$0.12 0.7208$\pm$0.12 0.6867$\pm$0.12 0.6650$\pm$0.12 0.6175$\pm$0.04 0.6867$\pm$0.12 0.6950$\pm$0.12 0.7242 $\pm$0.08 0.6583$\pm$ 0.00 0.7458$\pm$0.08
+CS 0.5462$\pm$0.03 0.5488$\pm$0.06 0.5788$\pm$0.07 0.5875$\pm$0.04 0.5825$\pm$0.05 0.6350$\pm$0.03 0.6075$\pm$0.02 0.5837$\pm$0.05 0.5525$\pm$0.04 0.5863$\pm$0.04 0.5000$\pm$ 0.00 0.6313$\pm$0.04
+HR 0.3758$\pm$0.02 0.3886$\pm$0.03 0.3917$\pm$0.02 0.3924$\pm$0.04 0.3773$\pm$0.02 0.3697$\pm$0.04 0.3333$\pm$0.00 0.4000$\pm$0.03 0.3758$\pm$0.02 0.3939$\pm$0.03 0.4167$\pm$ 0.00 0.4326$\pm$0.02
+ZO 0.6624$\pm$0.12 0.6733$\pm$0.11 0.6743$\pm$0.11 0.6752$\pm$0.10 0.6693$\pm$0.10 0.6802$\pm$0.08 0.7059$\pm$0.05 0.6733$\pm$0.10 0.6743$\pm$0.11 0.6228$\pm$0.07 0.6832$\pm$ 0.00 0.7792$\pm$0.10
+BC 0.5472$\pm$0.04 0.5594$\pm$0.07 0.6392$\pm$0.09 0.6346$\pm$0.09 0.5413$\pm$0.06 0.6140$\pm$0.09 0.5266$\pm$0.06 0.6052$\pm$0.10 0.6185$\pm$0.10 0.5804$\pm$0.08 0.6399$\pm$ 0.00 0.6650$\pm$0.05
+LG 0.4277$\pm$0.06 0.4446$\pm$0.04 0.4547$\pm$0.06 0.4905$\pm$0.06 0.4466$\pm$0.05 0.5095$\pm$0.08 0.4959$\pm$0.05 0.4959$\pm$0.05 0.5101$\pm$0.04 0.4743$\pm$0.04 0.5308$\pm$ 0.00 0.5426$\pm$0.02
+TT 0.5569$\pm$0.05 0.5593$\pm$0.04 0.5563$\pm$0.04 0.5408$\pm$0.03 0.5386$\pm$0.03 0.5637$\pm$0.03 0.5640$\pm$0.04 0.5465$\pm$0.04 0.5471$\pm$0.04 0.6078$\pm$ 0.03 0.5021$\pm$ 0.00 0.5785$\pm$0.03
+AC 0.7281$\pm$0.12 0.7783$\pm$0.08 0.6335$\pm$0.12 0.6354$\pm$0.12 0.7683$\pm$0.08 0.7488$\pm$0.14 0.7703$\pm$0.08 0.6762$\pm$0.14 0.6293$\pm$0.14 0.7717$\pm$0.10 0.5565$\pm$0.00 0.8206$\pm$0.08
+VT 0.8625$\pm$0.01 0.8699$\pm$0.00 0.8699$\pm$0.00 0.8761$\pm$0.00 0.8667$\pm$0.00 0.8531$\pm$0.09 0.8736$\pm$0.00 0.8736$\pm$0.00 0.8713$\pm$0.00 0.8667$\pm$0.00 0.6092$\pm$0.00 0.8943$\pm$0.00
+$\overline{\textit{AR}}$ 8.75 7.38 6.54 5.83 8.04 5.33 6.08 6.29 7.17 7.2083 8.2917 1.08
+Table 2: Clustering performance evaluated by ARI (with range $[-1,1]$). “$\overline{\textit{AR}}$” row reports the average performance rankings.
+Datasets KMD LSM JDM CBDM UDMC DLC HARR HDC ADC MCDC AMPHM OCL
+SB 0.7638$\pm$0.18 0.7555$\pm$0.19 0.7232$\pm$0.19 0.7407$\pm$0.19 0.7307$\pm$0.22 0.7148$\pm$0.20 0.8650$\pm$0.22 0.7350$\pm$0.19 0.6968$\pm$0.17 0.2736$\pm$0.18 0.5949$\pm$0.00 0.9620$\pm$0.12
+NS 0.0376$\pm$0.03 0.0341$\pm$0.02 0.0283$\pm$0.02 -- 0.0393$\pm$0.03 0.0779$\pm$0.08 0.0309$\pm$0.03 0.0376$\pm$0.03 0.0376$\pm$0.03 0.0337$\pm$0.02 -- 0.1015$\pm$0.06
+AP 0.0022$\pm$0.01 0.0171$\pm$0.02 0.0530$\pm$0.03 0.0322$\pm$0.02 0.0178$\pm$0.02 0.0355$\pm$0.00 0.0076$\pm$0.02 0.0150$\pm$0.02 0.0066$\pm$0.02 0.0005$\pm$0.02 -0.0055$\pm$0.00 0.0524$\pm$0.00
+DS 0.1077$\pm$0.13 0.1077$\pm$0.13 0.2347$\pm$0.20 0.2347$\pm$0.20 0.1844$\pm$0.20 0.1556$\pm$0.19 0.0512$\pm$0.05 0.1844$\pm$0.20 0.1921$\pm$0.19 0.2108$\pm$0.14 0.0923$\pm$0.00 0.2590$\pm$0.16
+CS -0.0023$\pm$0.01 0.0088$\pm$0.04 0.0295$\pm$0.04 0.0250$\pm$0.03 0.0248$\pm$0.04 0.0647$\pm$0.03 0.0342$\pm$0.02 0.0234$\pm$0.03 0.0040$\pm$0.02 0.0187$\pm$0.03 -0.0142$\pm$0.00 0.0627$\pm$0.03
+HR -0.0081$\pm$0.01 -0.0008$\pm$0.01 -0.0013$\pm$0.01 -0.0010$\pm$0.01 -0.0046$\pm$0.01 -0.0033$\pm$0.01 -0.0149$\pm$0.00 0.0038$\pm$0.02 -0.0064$\pm$0.01 -0.0010$\pm$0.04 0.0094$\pm$0.00 0.0368$\pm$0.01
+ZO 0.5787$\pm$0.17 0.6085$\pm$0.17 0.5996$\pm$0.17 0.6091$\pm$0.15 0.6063$\pm$0.16 0.6094$\pm$0.12 0.6442$\pm$0.07 0.6138$\pm$0.15 0.6143$\pm$0.15 0.4760$\pm$0.16 0.7515$\pm$0.00 0.7536$\pm$0.11
+BC 0.0046$\pm$0.02 0.0240$\pm$0.05 0.0811$\pm$0.07 0.0777$\pm$0.07 0.0128$\pm$0.04 0.0523$\pm$0.05 -0.0033$\pm$0.00 0.0605$\pm$0.08 0.0692$\pm$0.08 0.0084$\pm$0.05 0.0393$\pm$0.00 0.0799$\pm$0.03
+LG 0.0843$\pm$0.05 0.0951$\pm$0.03 0.1027$\pm$0.06 0.1432$\pm$0.07 0.1018$\pm$0.04 0.1544$\pm$0.09 0.1715$\pm$0.04 0.1095$\pm$0.06 0.1262$\pm$0.05 -0.0012$\pm$0.02 0.1293$\pm$0.00 0.1552$\pm$0.03
+TT 0.0170$\pm$0.03 0.0184$\pm$0.02 0.0181$\pm$0.02 0.0095$\pm$0.01 0.0091$\pm$0.01 0.0156$\pm$0.01 0.0198$\pm$0.02 0.0120$\pm$0.02 0.0124$\pm$0.02 0.0051$\pm$0.03 -0.0022$\pm$0.00 0.0226$\pm$0.02
+AC 0.2596$\pm$0.18 0.3336$\pm$0.12 0.1210$\pm$0.15 0.1199$\pm$0.13 0.3092$\pm$0.11 0.3184$\pm$0.22 0.3107$\pm$0.11 0.1913$\pm$0.18 0.1349$\pm$0.18 0.3263$\pm$0.19 0.0011$\pm$0.00 0.4313$\pm$0.16
+VT 0.5247$\pm$0.02 0.5463$\pm$0.01 0.5463$\pm$0.01 0.5648$\pm$0.01 0.5368$\pm$0.01 0.5200$\pm$0.18 0.5572$\pm$0.00 0.5572$\pm$0.00 0.5503$\pm$0.00 0.5367$\pm$0.00 0.0013$\pm$0.00 0.6207$\pm$0.00
+$\overline{\textit{AR}}$ 8.71 6.42 6.00 5.83 7.29 5.58 6.04 5.92 7.00 8.75 9.125 1.33
+[…]
+The Appendix presents supplementary experimental settings and results including: (1) Detailed dataset statistics (A.3); (2) Significant tests (A.4); (3) Performance evaluated by NMI (A.5); (4) Ablation studies evaluated by ARI and NMI (A.6); (5) Additive efficiency evaluation (A.7); (6) Case study (A.8); (7) Complete convergence evaluation (A.9); (8) Performance on mixed datasets (A.10); (9) Performance evaluation of applying OCL to Numerical Data (A.11); (10) Performance on large-scale datasets (A.12).
+4.2 Clustering Performance Evaluation
+The clustering results obtained by different algorithms on various datasets in terms of different validity indices are shown in Tables 1 and 2.
+The best and second-best results on each dataset are highlighted in bold and underlined, respectively.
+The observations include the following four aspects: (1) Overall, OCL reaches the best result on most datasets, indicating its superiority in clustering.
+(2) On the CA index, although OCL does not perform significantly better than the second-best method on the AP and LG datasets, the second-best methods are different on these datasets, but OCL has stable performance on them.
+(3) On the CA index, although OCL does not perform the best on CS and TT datasets, its performance is close to the best result (the difference is less than 0.005 in CS dataset, less than 0.03 in TT dataset), which still proves OCL’s competitiveness.
+(4) In terms of the ARI index, there are four datasets on which OCL does not perform the best, namely AP, CS, BC, and LG.
+However, similarly, the differences between OCL and the best-performing algorithms on AP, CS, and BC datasets are very small.
+In addition, the results of CBDM and AMPHM are not reported on the NS dataset because all the NS’s attributes are independent of each other, making the inter-attribute dependence-based methods like CBDM and AMPHM fail in measuring distances.
+Table 3: CA performance of ablated OCL variants.
+The green arrows indicate the expected performance improvements achieved over a more basic variant.
+OD, OL, OI, and $\overline{\textit{AR}}$ indicate order distance, order learning, order information, and average rank, respectively.
+OCL Components Mark Datasets $\overline{\textit{AR}}$
+OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
+$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9830$\uparrow$   0.3573$\uparrow$   0.6222$\uparrow$   0.7458$\uparrow$   0.6313   0.4326$\uparrow$   0.7792$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   0.5785$\uparrow$   0.8206$\uparrow$   0.8943$\uparrow$   1.17$\uparrow$
+$\checkmark$ $\checkmark$ I   0.9149$\uparrow$   0.3364$\uparrow$   0.6085$\uparrow$   0.7283$\uparrow$   0.6500$\uparrow$   0.3765$\uparrow$   0.7525$\uparrow$   0.5643$\uparrow$   0.4872$\uparrow$   0.5784$\uparrow$   0.7546$\uparrow$   0.8942$\uparrow$   2.00$\uparrow$
+$\checkmark$ II   0.8681$\uparrow$   0.3359$\uparrow$   0.6069$\uparrow$ 0.
+7125$\uparrow$   0.6388$\uparrow$ 0.3712   0.7000$\uparrow$   0.5636$\uparrow$   0.4453$\uparrow$   0.5762$\uparrow$   0.7545$\uparrow$   0.8828$\uparrow$   2.92$\uparrow$
+III 0.8468 0.3208 0.5392 0.6525 0.5463 0.3758 0.6624 0.5472 0.4277 0.5569 0.7281 0.8625 3.92
+4.3 Ablation Studies
+To demonstrate the effectiveness of the core components of OCL and to verify our hypothesis on the orders of categorical data attributes, several ablated variants of OCL are compared from both algorithm and data perspectives.
+From the algorithm aspect, to evaluate the effectiveness of the proposed Order Distance (OD), we compare OCL with its variant OCL-I, which uses normalized equidistant order distance without considering the probability distribution of possible values.
+To evaluate the effectiveness of the Order Learning (OL) mechanism, we further remove the process of order learning from OCL-I and only update the order once in the first iteration, which forms OCL-II.
+To completely discard Order Information (OI), we further make OCL-II use only the traditional Hamming distance, thus forming OCL-III.
+Table 3 compares the three OCL variants.
+It can be observed that the performance of OCL is superior to all variants.
+More specifically, OCL outperforms OCL-I significantly on 11 out of all 12 datasets, indicating that OD can effectively utilize the distribution information of the attribute values within different clusters for reasonable order distance measurement.
+Furthermore, the performance of OCL-I is also significantly better than OCL-II on all 12 datasets, which verifies the effectiveness of the OL mechanism in searching optimal orders during clustering.
+In addition, compared to OCL-III that completely disregards OI, OCL-II has better performance on 11 out of all 12 datasets, which proves that even a roughly estimated initial order can still effectively improve clustering performance.
+Table 4: CA performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
+Learn Order for Mark Datasets $\overline{\textit{AR}}$
+OA NA NS AP CS HR BC LG
+OCL   0.3573$\uparrow$   0.6222$\uparrow$ 0.6313   0.4326$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   1.17$\uparrow$
+LNRO 0.3281   0.6074$\uparrow$   0.6350$\uparrow$   0.3985$\uparrow$   0.5476$\uparrow$   0.4818$\uparrow$   2.00$\uparrow$
+RNRO 0.3308 0.5984 0.6225 0.3902 0.5276 0.4493 2.83
+From the data aspect, to validate our hypothesis that correct order information is crucial to categorical data clustering, we conduct ablation experiments on datasets containing both nominal and ordinal attributes.
+We compare the proposed OCL, which learns orders for both the nominal and ordinal attributes, to two different variants: (1) LNRO (Learn Nominal Reserve Ordinal), which retains the semantic order information of the original ordinal attribute and only learns orders for the nominal attributes, and (2) RNRO (Reserve Nominal Reserve Ordinal), which discards the order learning process for both nominal and ordinal attributes.
+Table 4 compares the clustering performance of the three variants.
+It can be observed that on five out of the six datasets, the performance of LNRO is superior to RNRO.
+This verifies our hypothesis that the nominal attributes also contain potential order information that is effective for clustering.
+We can also find that the OCL outperforms LNRO on five out of the six datasets.
+This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering than the original semantic orders.
+Figure 2: Convergence curves of OCL on the three datasets.
+4.4 Convergence and Efficiency Evaluation
+To evaluate the convergence of OCL, we execute it on three datasets i.e., SB, NS, and AP.
+The convergence curves are presented in Figure 2.
+The horizontal and vertical axes represent the number of learning iterations and the value of the objective function $L$, respectively.
+The blue triangles mark the iterations of order updates, while the red square indicates the iteration that OCL converges.
+It can be observed that, after each order update, the value of $L$ further decreases, indicating that the proposed order learning mechanism is consistent with the optimization of the clustering objective.
+Moreover, OCL always completes the learning process within 30 iterations with at most three updates of the orders.
+This reflects that OCL converges efficiently with guarantee, conforms to the convergence analysis in Section 3.5.
+Figure 3: Execution time (y-axis) on synthetic datasets with different numbers of samples $n$ and attributes $d$ (x-axis).
+Furthermore, to evaluate the execution efficiency of OCL, we construct large synthetic datasets with different numbers of samples and attributes to investigate the execution time of OCL in comparison with other counterparts.
+The results reported in Figure 3 show that the execution time of OCL is better than UDMC, AMPHM, and DLC on large-scale datasets, and better than HARR, UDMC, AMPHM, and CBDM on high-dimensional datasets.
+Additionally, as the number of samples or attributes increases, the execution time of OCL increases linearly, which is consistent with the time complexity analysis in Section 3.5.
+In summary, OCL is efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
+Figure 4: t-SNE visual comparison of CBDM (left), ADC (middle), and OCL (right) on the VT dataset.
+4.5 Visualization of Clustering Effect
+To demonstrate the cluster discrimination ability of the OCL method, the attribute distances learned by OCL, CBDM, and ADC are utilized to encode the attributes of the VT dataset for 2-dimensional t-SNE (Maaten & Hinton, 2008) visualization.
+It can be seen from Figure 4 that OCL has significantly better cluster discrimination ability, as it performs clustering task-oriented distance learning to better suit the exploration of the $k$ clusters of the VT dataset.
+5 Concluding Remarks
+This paper introduces and validates a new finding that knowing the correct order relation of attribute values is critical to cluster analysis of categorical data.
+Accordingly, the OCL paradigm is proposed allowing joint learning of data partitions and order relation-guided distances.
+Its learning mechanism can infer a more appropriate order among attribute values based on the current clusters, thus yielding more compact clusters based on the updated order.
+The data partition and order learning are iteratively executed to approach the optimal clustering solution.
+As a result, more accurate clustering results and insightful orders can be obtained.
+OCL converges quickly with guarantee, and extensive experiments on 12 real benchmark datasets fully demonstrate its superiority in comparison with state-of-the-art clustering methods.
+A case study has also been provided to show the rationality of the learned order in understanding categorical data clusters.
+The proposed OCL is also not exempt from limitations.
+For instance, the natural connection between the treatment of categorical and numerical attributes remains to be explored.
+It would also be promising to extend OCL to more complex scenarios, e.g., non-stationary mixed dataset with an unknown number of imbalanced clusters.
+Appendix A Appendix
+A.1 Frequently Used Symbols and Detailed Problem Formulation with Examples
+Frequently used symbols and explanations are provided in Table 5.
+To ensure a more intuitive understanding of the problem formulation, we also provide it with more detailed explanations and specific dataset examples below.
+The problem of categorical data clustering with order learning is formulated as follows.
+Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples (e.g., a collection of $n$ clients of a bank).
+Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$ (e.g., occupation, gender, priority, etc.).
+Each attribute $\mathbf{a}_{r}$ (e.g., occupation) can be denoted as a column vector $\mathbf{a}_{r}=[x_{1,r},x_{2,r},...,x_{n,r}]$ composed of the occupation description of all the $n$ clients.
+For the categorical attribute “occupation” $\mathbf{a}_{r}$, all its $n$ values are taken from a limited number of qualitative possible values $V_{r}=\{v_{r,1},v_{r,2},...,v_{r,l_{r}}\}$, e.g., $\{\text{lawyer},\text{doctor},...,\text{scientist}\}$.
+The subscripts indicate the sequential number of an attribute and the sequential number of the attribute’s possible value.For example, doctor is the second value in the $r$-th attribute “occupation”, so we have $v_{r,2}=\text{doctor}$.
+We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data.
+Please note that the sequential numbers of possible values only distinguish them, and do not indicate their order.
+Table 5: Frequently used symbols. Note that we uniformly use lowercase, uppercase, bold lowercase, and bold uppercase to indicate value, set, vector, and matrix, respectively.
+Symbol Explanation
+$X$, $A$, $O$, and $C$ Dataset, attribute set, order set, and cluster set
+$\mathbf{x}_{i}$ and $x_{i,r}$ $i$-th data object and $r$-th value of $\mathbf{x}_{i}$
+$\mathbf{a}_{r}$, $V_{r}$, and $v_{r,g}$ $r$-th attribute, its value set, and $g$-th value of $V_{r}$
+$l_{r}$ Total number of possible values in $V_{r}$
+$O_{r}$ Order values (integers) of $\mathbf{a}_{r}$’s possible values $V_{r}$
+$o(v_{r,g})$ Order value (integer) of possible value $v_{r,g}$
+$\mathbf{Q}$ and $C_{m}$ Object-cluster affiliation matrix and $m$-th cluster
+$q_{i,m}$ $(i,m)$-th entry of $\mathbf{Q}$ indicating affiliation of $\mathbf{x}_{i}$ to $C_{m}$
+$\Theta(\mathbf{x}_{i},C_{m};O)$ Overall order distance between $\mathbf{x}_{i}$ and $C_{m}$
+$\theta(x_{i,r},\textbf{p}_{m,r})$ Order distance between $\mathbf{x}_{i}$ and $C_{m}$ reflected by $\mathbf{a}_{r}$
+$\mathbf{D}$ and $\mathbf{d}_{i,r}$ Order distance matrix and its $(i,r)$-th entry
+$d_{i,r,g}$ Order distance between $x_{i,r}$ and $v_{r,g}$, also the $g$-th value of $\mathbf{d}_{i,r}$
+$\mathbf{P}$ and $\mathbf{p}_{m,r}$ Conditional probability matrix and its $(m,r)$-th entry
+$p_{m,r,g}$ Occurrence probability of $v_{r,g}$ in $C_{m}$, also the $g$-th value of $\mathbf{p}_{m,r}$
+$O_{r,m}$ Order values of $\mathbf{a}_{r}$’s possible values $V_{r}$ obtained through $C_{m}$
+$L_{r,m}$ A fraction of $L$ jointly contributed by $\mathbf{a}_{r}$ and $C_{m}$
+Assume that there is an optimal order indicating an appropriate underlying distance structure of possible values w.r.t.
+a clustering task.
+We denote the optimal order of an attribute $\mathbf{a}_{r}$ as $O^{*}_{r}=\{o^{*}(v_{r,1}),o^{*}(v_{r,2}),...,o^{*}(v_{r,l_{r}})\}$ where the superscript “*” marks the optimum and the value of $o^{*}(v_{r,g})$ is an integer reflecting the unique ranking of $v_{r,g}$ among all the $l_{r}$ possible values in $V_{r}$.
+By still taking the “occupation” attribute as an example, if its three possible values $v_{r,1}=\text{lawyer}$, $v_{r,2}=\text{doctor}$, and $v_{r,3}=\text{scientist}$, ranking second (i.e., $o(\text{lawyer})=2$), third (i.e., $o(\text{doctor})=3$), and first (i.e., $o(\text{scientist})=1$), respectively.
+Then the corresponding order of these values are $O_{r}=\{o(\text{lawyer}),o(\text{doctor}),o(\text{scientist})\}=\{2,3,1\}$.
+Such order reflects a distance structure that roughly satisfies $\text{dist}(v_{r,g},v_{r,h})\propto|o(v_{r,g})-o(v_{r,h})|$.
+More specifically, the following distance relationship should hold:
+1. dist(doctor, scientist) $>$ dist(doctor, lawyer),
+2. dist(doctor, scientist) $>$ dist(lawyer,scientist),
+because “doctor” and “scientist” are further away in order, compared to “lawyer” and “doctor” that are adjacent in order.
+We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$.
+Then the order information reflecting the distance structure can guide the partition of data objects during clustering.
+Our approach, i.e., learning order relations, is more generative than most existing approaches that directly define the distance between attribute values.
+We focus more on how the intrinsic relationship between values affects the distance between samples in forming clusters, rather than simply defining the clustering-irrelevant distance between values based on data statistics, e.g., occurrence probability of values in an attribute.
+For a crisp partitional clustering task to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
+$L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$ (15)
+where $\mathbf{Q}$ is an $n\times k$ matrix with its $(i,m)$-th entry $q_{i,m}$ indicating the affiliation between sample $\mathbf{x}_{i}$ and cluster $C_{m}$.
+Specifically, $q_{i,m}=1$ indicates that $\mathbf{x}_{i}$ belongs to $C_{m}$, while $q_{i,m}=0$ indicates that $\mathbf{x}_{i}$ belongs to a cluster other than $C_{m}$, which can be written as:
+$q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$ (16)
+with $\sum_{i=m}^{k}q_{i,m}=1$.
+$\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$.
+Therefore, the problem can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$.
+When the algorithm converges, a sub-optimal solution with $k$ compact bank client clusters is expected to be obtained.
+In general, “compact” means that the bank clients with more similar characteristics (i.e., with more similar value descriptions across all the attributes) are gathered into the same cluster, and different clusters are relatively distinguishable.
+A.2 Detailed Complexity and Convergence Analysis
+We analyze the time complexity to show the scalability of OCL, and then provide its convergence analysis.
+Theorem A.1.
 The time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$.
@@
 □
-Remark 3.
+Remark A.2.
 Scalability of OCL.
-The essential characteristic of categorical data is that the number of possible values for an attribute is finite and small, usually ranging from 2 to 8 for real benchmark datasets as shown in the Appendix.
+The essential characteristic of categorical data is that the number of possible values for an attribute is finite and small, usually ranging from 2 to 8 for real benchmark datasets as shown in Table 6 of the Appendix.
 Accordingly, $\varsigma$ can be treated as a small constant, and the time complexity is simplified to $O(EInkd)$, which is linear to the scaling of $n$ or $d$.
-Theorem 2.
+Theorem A.3.
 OCL algorithm converges to a local minimum in a finite number of iterations.
 Proof. We prove the convergence of the inner and outer loop, i.e., lines 4 - 9 and lines 2 - 13 of Algorithm 1, respectively.
-Lemma 0.
+Lemma A.4.
 The inner loop converges to a local minimum of $L(\mathbf{Q},\hat{O})$ in a finite number of iterations with fixed $\hat{O}$.
@@
 Hence, the result follows.
-Lemma 0.
+Lemma A.5.
 The outer loop converges to a local minimum of $L(\mathbf{Q},O)$ in a finite number of iterations.
@@
 Since both the inner loop and outer loop of Algorithm 1 converge to a local minimum, convergence of OCL can be confirmed. □
-4. Experiments
-Different experiments have been designed to evaluate the performance of the proposed order learning strategy as well as the corresponding clustering algorithm.
-Detailed experimental settings and the obtained results are introduced as follows.
-Table 2.
-Statistics of the 12 datasets.
-$d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, and $n$ are the numbers of nominal attributes, ordinal attributes, and samples, respectively.
-$k^{*}$ is the number of true clusters.
-No. Dataset Abbrev. $d^{\text{\{nom\}}}$ $d^{\text{\{ord\}}}$ $n$ $k^{*}$
-1 Soybean (small) SB 35 0 47 4
-2 Nursery School NS 1 7 12960 4
-3 Amphibians AP 6 6 189 2
-4 Inflammations Diagnosis DS 0 5 120 2
-5 Caesarian Section CS 1 3 80 2
-6 Hayes-Roth HR 2 2 132 3
-7 Zoo ZO 15 0 101 7
-8 Breast Cancer BC 5 4 286 2
-9 Lymphography LG 15 3 148 4
-10 Tic-Tac-Toe TT 9 0 958 2
-11 Australia Credit AC 0 8 690 2
-12 Congressional Voting VT 16 0 435 2
-4.1. Experimental Settings
-Four types of experiments have been conducted to evaluate the proposed method.
-A brief outline of the experiments is given below: (1) Clustering performance evaluation with significance test (Section 4.2): We compared the proposed OCL with conventional and state-of-the-art methods on different types of datasets.
-To statistically analyze the performance differences between variant approaches, we also conducted a significance test on the clustering results; (2) Ablation studies (Section 4.3): To more specifically demonstrate the effectiveness of the core components of OCL and investigate the role of learned order in improving clustering performance, we conduct ablation studies from the algorithm aspect and data aspect.
-(3) Convergence and efficiency evaluation (Section 4.4): The values of the objective function after each learning iteration have been recorded to study the convergence of OCL.
-Its efficiency has also been studied with experiments on large datasets with scaling sizes of samples $n$ and attributes $d$.
-(4) Visualization of clustering effect (Section 4.5): To visually demonstrate the effectiveness of the order distance metric learned by OCL, we use it and the counterparts to numerically encode the data, and utilize the t-SNE for visualization.
-Nine counterparts have been selected for the comparative study.
-The counterparts include traditional K-MoDes (KMD) algorithm (Huang, 1998), popular entropy-based metric LSM (Lin, 1998), and context-based distance measurement CBDM (Ienco et al., 2009).
-The latter two are representative measures proposed for categorical data, both of which can be embedded into KMD to conduct clustering.
-Besides, some state-of-the-art methods have also been taken into account, including Jia’s Distance Metric (JDM) (Jia et al., 2016), Unified Distance Metric based Clustering (UDMC) (Zhang and Cheung, 2022b), HD based Clustering (HDC) (Zhang and Cheung, 2022a), projection based clustering Het2Hom (H2H) (Zhang et al., 2022), Distance Learning based Clustering (DLC) algorithm (Zhang and Cheung, 2020), and graph metric based clustering algorithm ADC (Zhang and Cheung, 2023).
-Please note that we have set the parameters (if any) of these methods to the values recommended in the corresponding paper.
-In the experiments, each clustering result was recorded based on the average performance of 10 runs.
-Three validity indices, i.e., clustering accuracy (CA) (He et al., 2005), ARI coefficients (Rand, 1971)(Gates and Ahn, 2017), and standardized mutual information NMI (Estévez et al., 2009), have been selected to evaluate the clustering performance.
-Usually, larger values of these indices indicate better clustering performance.
-We also conduct the Bonferroni Dunn (BD) test [^3] The Friedman test (Friedman, 1937) is performed first to ensure a meaningful BD post-hoc test.
-(Demšar, 2006) on the performance of different methods and calculate the critical difference (CD) interval to statistically analyze the superiority of our OCL.
+A.3 Supplementary Experimental Settings
+A.3.1 Detailed settings of the demo experiment in Figure 1 of the Introduction
 […]
-Table 3. Clustering performance evaluated by CA. “$\overline{\textit{AR}}$” row reports the average performance rankings.
-Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
-SB 0.8468$\pm$0.15 0.8106$\pm$0.16 0.7894$\pm$0.14 0.7957$\pm$0.15 0.7915$\pm$0.18 0.8000$\pm$0.14 0.9234$\pm$0.12 0.7915$\pm$0.15 0.7660$\pm$0.13 0.9830$\pm$0.05
-NS 0.3208$\pm$0.05 0.2975$\pm$0.04 0.2980$\pm$0.04 - 0.3074$\pm$0.05 0.3386$\pm$0.07 0.3113$\pm$0.05 0.3208$\pm$0.05 0.3208$\pm$0.05 0.3573$\pm$0.05
-AP 0.5392$\pm$0.03 0.5656$\pm$0.04 0.6159$\pm$0.04 0.5926$\pm$0.03 0.5698$\pm$0.03 0.6032$\pm$0.00 0.5603$\pm$0.03 0.5608$\pm$0.04 0.5529$\pm$0.04 0.6222$\pm$0.00
-DS 0.6525$\pm$0.08 0.6525$\pm$0.08 0.7208$\pm$0.12 0.7208$\pm$0.12 0.6867$\pm$0.12 0.6650$\pm$0.12 0.6175$\pm$0.04 0.6867$\pm$0.12 0.6950$\pm$0.12 0.7458$\pm$0.08
-CS 0.5462$\pm$0.03 0.5488$\pm$0.06 0.5788$\pm$0.07 0.5875$\pm$0.04 0.5825$\pm$0.05 0.6350$\pm$0.03 0.6075$\pm$0.02 0.5837$\pm$0.05 0.5525$\pm$0.04 0.6313$\pm$0.04
-HR 0.3758$\pm$0.02 0.3886$\pm$0.03 0.3917$\pm$0.02 0.3924$\pm$0.04 0.3773$\pm$0.02 0.3697$\pm$0.04 0.3333$\pm$0.00 0.4000$\pm$0.03 0.3758$\pm$0.02 0.4326$\pm$0.02
-ZO 0.6624$\pm$0.12 0.6733$\pm$0.11 0.6743$\pm$0.11 0.6752$\pm$0.10 0.6693$\pm$0.10 0.6802$\pm$0.08 0.7059$\pm$0.05 0.6733$\pm$0.10 0.6743$\pm$0.11 0.7792$\pm$0.10
-BC 0.5472$\pm$0.04 0.5594$\pm$0.07 0.6392$\pm$0.09 0.6346$\pm$0.09 0.5413$\pm$0.06 0.6140$\pm$0.09 0.5266$\pm$0.06 0.6052$\pm$0.10 0.6185$\pm$0.10 0.6650$\pm$0.05
-LG 0.4277$\pm$0.06 0.4446$\pm$0.04 0.4547$\pm$0.06 0.4905$\pm$0.06 0.4466$\pm$0.05 0.5095$\pm$0.08 0.4959$\pm$0.05 0.4959$\pm$0.05 0.5101$\pm$0.04 0.5426$\pm$0.02
-TT 0.5569$\pm$0.05 0.5593$\pm$0.04 0.5563$\pm$0.04 0.5408$\pm$0.03 0.5386$\pm$0.03 0.5637$\pm$0.03 0.5640$\pm$0.04 0.5465$\pm$0.04 0.5471$\pm$0.04 0.5785$\pm$0.03
-AC 0.7281$\pm$0.12 0.7783$\pm$0.08 0.6335$\pm$0.12 0.6354$\pm$0.12 0.7683$\pm$0.08 0.7488$\pm$0.14 0.7703$\pm$0.08 0.6762$\pm$0.14 0.6293$\pm$0.14 0.8206$\pm$0.08
-VT 0.8625$\pm$0.01 0.8699$\pm$0.00 0.8699$\pm$0.00 0.8761$\pm$0.00 0.8667$\pm$0.00 0.8531$\pm$0.09 0.8736$\pm$0.00 0.8736$\pm$0.00 0.8713$\pm$0.00 0.8943$\pm$0.00
-$\overline{\textit{AR}}$ 7.58 6.46 5.71 5.13 7.08 4.67 5.33 5.63 6.33 1.08
-Table 4. Clustering performance evaluated by ARI (with range [-1,1]). “$\overline{\textit{AR}}$” row reports the average performance rankings.
-Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
-SB 0.7638$\pm$0.18 0.7555$\pm$0.19 0.7232$\pm$0.19 0.7407$\pm$0.19 0.7307$\pm$0.22 0.7148$\pm$0.20 0.8650$\pm$0.22 0.7350$\pm$0.19 0.6968$\pm$0.17 0.9620$\pm$0.12
-NS 0.0376$\pm$0.03 0.0341$\pm$0.02 0.0283$\pm$0.02 - 0.0393$\pm$0.03 0.0779$\pm$0.08 0.0309$\pm$0.03 0.0376$\pm$0.03 0.0376$\pm$0.03 0.1015$\pm$0.06
-AP 0.0022$\pm$0.01 0.0171$\pm$0.02 0.0530$\pm$0.03 0.0322$\pm$0.02 0.0178$\pm$0.02 0.0355$\pm$0.00 0.0076$\pm$0.02 0.0150$\pm$0.02 0.0066$\pm$0.02 0.0524$\pm$0.00
-DS 0.1077$\pm$0.13 0.1077$\pm$0.13 0.2347$\pm$0.20 0.2347$\pm$0.20 0.1844$\pm$0.20 0.1556$\pm$0.19 0.0512$\pm$0.05 0.1844$\pm$0.20 0.1921$\pm$0.19 0.2590$\pm$0.16
-CS -0.0023$\pm$0.01 0.0088$\pm$0.04 0.0295$\pm$0.04 0.0250$\pm$0.03 0.0248$\pm$0.04 0.0647$\pm$0.03 0.0342$\pm$0.02 0.0234$\pm$0.03 0.0040$\pm$0.02 0.0627$\pm$0.03
-HR -0.0081$\pm$0.01 -0.0008$\pm$0.01 -0.0013$\pm$0.01 -0.0010$\pm$0.01 -0.0046$\pm$0.01 -0.0033$\pm$0.01 -0.0149$\pm$0.00 0.0038$\pm$0.02 -0.0064$\pm$0.01 0.0368$\pm$0.01
-ZO 0.5787$\pm$0.17 0.6085$\pm$0.17 0.5996$\pm$0.17 0.6091$\pm$0.15 0.6063$\pm$0.16 0.6094$\pm$0.12 0.6442$\pm$0.07 0.6138$\pm$0.15 0.6143$\pm$0.15 0.7536$\pm$0.11
-BC 0.0046$\pm$0.02 0.0240$\pm$0.05 0.0811$\pm$0.07 0.0777$\pm$0.07 0.0128$\pm$0.04 0.0523$\pm$0.05 -0.0033$\pm$0.00 0.0605$\pm$0.08 0.0692$\pm$0.08 0.0799$\pm$0.03
-LG 0.0843$\pm$0.05 0.0951$\pm$0.03 0.1027$\pm$0.06 0.1432$\pm$0.07 0.1018$\pm$0.04 0.1544$\pm$0.09 0.1715$\pm$0.04 0.1095$\pm$0.06 0.1262$\pm$0.05 0.1552$\pm$0.03
-TT 0.0170$\pm$0.03 0.0184$\pm$0.02 0.0181$\pm$0.02 0.0095$\pm$0.01 0.0091$\pm$0.01 0.0156$\pm$0.01 0.0198$\pm$0.02 0.0120$\pm$0.02 0.0124$\pm$0.02 0.0226$\pm$0.02
-AC 0.2596$\pm$0.18 0.3336$\pm$0.12 0.1210$\pm$0.15 0.1199$\pm$0.13 0.3092$\pm$0.11 0.3184$\pm$0.22 0.3107$\pm$0.11 0.1913$\pm$0.18 0.1349$\pm$0.18 0.4313$\pm$0.16
-VT 0.5247$\pm$0.02 0.5463$\pm$0.01 0.5463$\pm$0.01 0.5648$\pm$0.01 0.5368$\pm$0.01 0.5200$\pm$0.18 0.5572$\pm$0.00 0.5572$\pm$0.00 0.5503$\pm$0.00 0.6207$\pm$0.00
-$\overline{\textit{AR}}$ 7.88 5.92 5.50 5.38 6.71 5.08 5.29 5.50 6.42 1.33
-4.2. Clustering Performance Evaluation
-In this section, we investigate the clustering performance of different algorithms and statistically analyze the superiority of OCL.
-4.2.1. Clustering performance on different datasets
-The clustering results obtained by different algorithms on various datasets in terms of different validity indices are shown in Tables 3 and 4.
-The best and second-best results on each dataset are highlighted in bold and underlined, respectively.
-The observations include the following four aspects: (1) Overall, OCL reaches the best result on most datasets, indicating its superiority in clustering.
-(2) On the CA index, although OCL does not perform significantly better than the second-best method on the AP and TT datasets, the second-best methods are different on these datasets, but OCL has stable performance on them.
-(3) On the CA index, although OCL does not perform the best on CS dataset, its performance is close to the best result (the difference is less than 0.005), which still proves OCL’s competitiveness.
-(4) In terms of the ARI index, there are four datasets on which OCL does not perform the best, namely AP, CS, BC, and LG.
-However, similarly, the differences between OCL and the best-performing algorithms on AP, CS, and BC datasets are very small.
-In addition, the result of CBDM is not reported on the NS dataset because all the NS’s attributes are independent of each other, making the inter-attribute dependence-based CBDM fail in measuring distances.
-4.2.2. Significance study
-We first implement the Friedman (Friedman, 1937) test on the average ranks reported in the last rows in Tables 3 and 4.
-The corresponding p-values are 0.000017 and 0.000091 respectively, both passing the test under the confidence of 99% (p-value $=0.01$).
-On this basis, results of the two-tailed BD test (Demšar, 2006) with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 2, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively.
+A.3.2 Detailed settings of the Experiments section
+To ensure a more comprehensive evaluation, we select 12 real datasets from different fields.
+All datasets are collected from the UCI machine learning repository (Dua & Karra Taniskidou, 2017) and their detailed statistical information is summarized in Table 6.
+All datasets are pre-processed by removing samples with missing values.
+14 attributes of the SB dataset are omitted since each of them has only one possible value, which is meaningless for clustering.
+Since we focus on categorical data clustering, the numerical attributes of the datasets AP, DS, and AC are also omitted, expecting for the performance evaluation of enhancing mixed data clustering in Appendix A.10.
+To evaluate the efficiency of OCL, large synthetic datasets are randomly generated with different attribute and sample scales.
+Specifically, we generate by (1) fixing the number of attributes at $d=20$ and increasing the number of samples $n$ from 10k to 100k with step-size 10k, and (2) fixing sample size at $n=2k$ and increasing the number of attributes $d$ from 1k to 10k with step-size 1k.
+Here we use ‘k’ to indicate ‘kilo’.
+Under these settings, the values of samples are randomly generated from five possible values on each attribute.
+The number of clusters $k$ is consistently set to five.
+Table 6: Detailed statistics of the 12 datasets.
+14 attributes of the SB dataset are omitted since each of them has only one possible value, which is meaningless for clustering.
+“$d^{\{nom\}}$”, “$d^{\{ord\}}$”, and “$d^{\{num\}}$” indicate the numbers of nominal, ordinal, and numerical attributes, respectively.
+“$l_{r}$” indicates the number of possible values of categorical (ordinal + nominal) attributes.
+“$mean(l_{r})$”, “$max(l_{r})$”, and “$min(l_{r})$” indicate the averaged, maximum, and minimum numbers of possible values for the categorical attributes of each dataset, respectively.
+“$n$” is the number of samples.
+$k^{*}$ is the true number of clusters.
+Data abbrev. $d^{\{nom\}}$ $d^{\{ord\}}$ $d^{\{num\}}$ $l_{r}$ $mean(l_{r})$ $max(l_{r})$ $min(l_{r})$ $n$ $k^{*}$
+Soybeans SB 35 0 0 7 2 3 3 2 4 4 2 2 3 2 2 4 4 2 2 2 2 2 2 27233244223224422222227\ 2\ 3\ 3\ 2\ 4\ 4\ 2\ 2\ 3\ 2\ 2\ 4\ 4\ 2\ 2\ 2\ 2\ 2\ 2\ 27 2 3 3 2 4 4 2 2 3 2 2 4 4 2 2 2 2 2 2 2 2.76 7 2 47 4
+Nursery School NS 1 7 0 3 5 4 4 3 2 3 3354432333\ 5\ 4\ 4\ 3\ 2\ 3\ 33 5 4 4 3 2 3 3 3.38 5 2 12960 4
+Amphibians AP 6 6 2 8 5 8 7 8 3 5 6 6 6 3 28587835666328\ 5\ 8\ 7\ 8\ 3\ 5\ 6\ 6\ 6\ 3\ 28 5 8 7 8 3 5 6 6 6 3 2 5.58 8 2 189 2
+Inflammations Diagnosis DS 0 5 1 2 2 2 2 $2$\ 2\ 2\ 22 2 2 2 2 2.00 2 2 120 2
+Caesarian Section CS 1 3 0 4 3 3 243324\ 3\ 3\ 24 3 3 2 3.00 4 2 80 2
+Hayes-Roth HR 2 2 0 3 4 4 $\ 4$\ 43 4 4 4 3.75 4 3 132 3
+Zoo ZO 15 0 0 2 2 2 2 2 2 2 2 2 2 2 2 6 2 2 $2$\ 2\ 2\ 2\ 2\ 2\ 2\ 2\ 2\ 2\ 2\ 6\ 2\ 2\ 22 2 2 2 2 2 2 2 2 2 2 2 6 2 2 2 2.25 6 2 101 7
+Breast Cancer BC 5 4 0 3 3 2 6 2 6 11 7 $3$\ 2\ 6\ 2\ 6\ 11\ 7\ 33 3 2 6 2 6 11 7 3 4.78 11 2 286 2
+Lym LG 15 3 0 3 4 8 4 2 2 2 2 2 2 2 3 4 4 8 3 2 23484222222234483223\ 4\ 8\ 4\ 2\ 2\ 2\ 2\ 2\ 2\ 2\ 3\ 4\ 4\ 8\ 3\ 2\ 23 4 8 4 2 2 2 2 2 2 2 3 4 4 8 3 2 2 3.28 8 2 148 4
+Tic-Tae-Toc TT 9 0 0 3 3 3 3 3 3 3 3 $3$\ 3\ 3\ 3\ 3\ 3\ 3\ 33 3 3 3 3 3 3 3 3 3.00 3 3 958 2
+Australia Credit AC 0 8 6 2 2 2 2 3 14 8 32222314832\ 2\ 2\ 2\ 3\ 14\ 8\ 32 2 2 2 3 14 8 3 4.50 14 2 690 2
+Congressional Voting VT 16 0 0 3 3 3 3 3 3 3 3 3 3 3 3 3 3 3 $3$\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 3\ 33 3 3 3 3 3 3 3 3 3 3 3 3 3 3 3 3.00 3 3 435 2
+A.4 Significance Study of Clustering Performance Evaluated by CA and ARI
+We first implement the Friedman (Friedman, 1937) test on the average ranks reported in the last rows in Tables 1 and 2.
+The corresponding p-values are 0.000035 and 0.000011 respectively, both passing the test under the confidence of 99% (p-value $=0.01$).
+On this basis, results of the two-tailed BD test (Demšar, 2006) for all counterparts except the two oldest comparison methods, LSM and KMD, with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 5, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively.
 According to the criterion presented in (Demšar, 2006), OCL is considered to perform significantly better than all the counterparts that fall outside the right boundary of the CD intervals.
 It can be seen that the proposed OCL is significantly superior to all counterparts.
-Figure 2. Results of the two-tailed BD tests w.r.t. the CA (a) and ARI (b) performance in Tables 3 and  4, respectively.
-4.3. Ablation Studies
-To demonstrate the effectiveness of the core components of OCL and to verify our hypothesis on the orders of categorical data attributes, we design several variants of OCL for ablation studies from both algorithm and data perspectives.
-From the algorithm aspect, to evaluate the effectiveness of the proposed Order Distance (OD), we compare OCL with its variant OCLI, which uses normalized equidistant order distance without considering the probability distribution of possible values.
-To evaluate the effectiveness of the Order Learning (OL) mechanism, we further remove the process of order learning from OCLI and only update the order once in the first iteration, which forms another variant OCLII.
-To completely discard Order Information (OI), we further make OCLII use only the traditional Hamming distance, thus forming the third variant OCLIII.
-Table 5 compares the clustering performance of OCL and the three variants.
-It can be observed that the performance of OCL is superior to all variants.
-More specifically, OCL outperforms OCLI significantly on 11 out of all 12 datasets.
-This indicates that OD can effectively utilize the distribution information of the attribute values within different clusters for reasonable order distance measurement.
-Furthermore, the performance of OCLI is also significantly better than OCLII on all 12 datasets, which verifies the effectiveness of the OL mechanism in searching optimal orders during clustering.
-In addition, compared to the OCLIII variant that completely disregards OI, OCLII has better performance on 11 out of all 12 datasets, which proves that even a roughly estimated initial order can still significantly and extensively improve clustering performance.
-Table 5.
-CA performance of ablated OCL variants formed by successively ablating the three core components, i.e., Order Distance (OD), Order Learning (OL), and Order Information (OI).
-The green arrows indicate the performance improvements of an OCL variant over its previous more basic variant.
-“$\overline{\textit{AR}}$” column reports the average performance rankings.
-OCL Components Variants Datasets $\overline{\textit{AR}}$
+Figure 5: Results of the two-tailed BD tests w.r.t. the CA (a) and ARI (b) performance in Tables 1 and  2, respectively.
+A.5 Clustering Performance Evaluated by NMI
+The clustering performance of different methods is compared in Table 7 evaluated by the NMI validity index.
+The best and second-best results on each dataset are highlighted in bold and underlined, respectively.
+Overall, OCL performs the best on 8 out of 12 datasets, indicating its superiority in general.
+Although OCL does not have the best NMI performance on CS, BC, LG, and TT datasets, it is not surpassed by much by the winners and still being very competitive.
+In addition, the result of CBDM on the NS dataset is not reported because the attributes of this dataset are independent of each other, making CBDM fail in measuring distances based on the interdependence of attributes.
+The corresponding significance test results w.r.t.
+NMI is demonstrated in Figure 6.
+We first conducted the Friedman test (Friedman, 1937) to demonstrate the significant differences in performance results, followed by the Bonferroni Dunn (BD) test (Demšar, 2006) to compare the performance of different methods and calculated the critical difference (CD) interval to statistically analyze the superiority of our OCL.
+The test is conducted by first implementing the Friedman test on the average performance ranks reported in the last rows in Table 7.
+The corresponding $p$-value is 0.00013, passing the Friedman test under $99\%$ confidence interval.
+On this basis, results of the two-tailed BD test (Demšar, 2006) for all counterparts except the two oldest comparison methods, LSM and KMD, with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 6, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively.
+All the compared methods fall outside the right boundary of the 90% CD intervals centered at OCL, indicating that OCL significantly outperforms them under the confidence interval of $90\%$.
+Table 7: Clustering performance evaluated by NMI (with range [0,1]). “$\overline{\textit{AR}}$” row reports the average performance rankings.
+Data KMD LSM JDM CBDM UDMC DLC HARR HDC ADC MCDC AMPHM OCL
+SB 0.8460$\pm$0.12 0.8418$\pm$0.13 0.8268$\pm$0.13 0.8393$\pm$0.12 0.8456$\pm$0.12 0.8298$\pm$0.12 0.9157$\pm$0.14 0.8513$\pm$0.11 0.8317$\pm$0.10 0.5344$\pm$0.17 0.7892$\pm$0.00 0.9757$\pm$0.08
+NS 0.0458$\pm$0.03 0.0441$\pm$0.02 0.0412$\pm$0.02 -- 0.0492$\pm$0.03 0.1348$\pm$0.10 0.0531$\pm$0.05 0.0458$\pm$0.03 0.0458$\pm$0.03 0.0504$\pm$0.03 -- 0.1492$\pm$0.10
+AP 0.0147$\pm$0.01 0.0321$\pm$0.03 0.0746$\pm$0.03 0.0469$\pm$0.03 0.0276$\pm$0.02 0.0649$\pm$0.00 0.0185$\pm$0.03 0.0384$\pm$0.04 0.0131$\pm$0.01 0.0194$\pm$0.02 0.0004$\pm$0.00 0.0955$\pm$0.01
+DS 0.1352$\pm$0.14 0.1352$\pm$0.14 0.2552$\pm$0.21 0.2552$\pm$0.21 0.2034$\pm$0.20 0.1682$\pm$0.21 0.0623$\pm$0.04 0.2034$\pm$0.20 0.2082$\pm$0.20 0.3083$\pm$0.13 0.0646$\pm$0.00 0.2732$\pm$0.18
+CS 0.0157$\pm$0.03 0.0221$\pm$0.04 0.0426$\pm$0.05 0.0353$\pm$0.03 0.0413$\pm$0.05 0.0840$\pm$0.03 0.0344$\pm$0.01 0.0346$\pm$0.03 0.0155$\pm$0.02 0.0310$\pm$0.03 0.0034$\pm$0.00 0.0808$\pm$0.04
+HR 0.0088$\pm$0.01 0.0144$\pm$0.01 0.0133$\pm$0.01 0.0110$\pm$0.01 0.0091$\pm$0.01 0.0100$\pm$0.01 0.0000$\pm$0.00 0.0182$\pm$0.02 0.0080$\pm$0.01 0.0407$\pm$0.08 0.0304$\pm$0.00 0.0456$\pm$0.02
+ZO 0.7172$\pm$0.08 0.7438$\pm$0.09 0.7395$\pm$0.09 0.7439$\pm$0.08 0.7541$\pm$0.09 0.7799$\pm$0.06 0.8324$\pm$0.04 0.7620$\pm$0.07 0.7731$\pm$0.08 0.7057$\pm$0.10 0.7190$\pm$0.00 0.8332$\pm$0.04
+BC 0.0044$\pm$0.01 0.0123$\pm$0.02 0.0355$\pm$0.03 0.0356$\pm$0.03 0.0068$\pm$0.01 0.0195$\pm$0.01 0.0017$\pm$0.00 0.0282$\pm$0.04 0.0318$\pm$0.03 0.0156$\pm$0.02 0.0091$\pm$0.00 0.0281$\pm$0.01
+LG 0.1309$\pm$0.04 0.1490$\pm$0.04 0.1477$\pm$0.05 0.1885$\pm$0.07 0.1537$\pm$0.04 0.1921$\pm$0.07 0.2047$\pm$0.04 0.1620$\pm$0.05 0.1680$\pm$0.03 0.0832$\pm$0.06 0.3300$\pm$0.00 0.1919$\pm$0.04
+TT 0.0105$\pm$0.01 0.0148$\pm$0.01 0.0146$\pm$0.02 0.0084$\pm$0.01 0.0074$\pm$0.01 0.0078$\pm$0.00 0.0163$\pm$0.02 0.0092$\pm$0.01 0.0094$\pm$0.01 0.0075$\pm$0.01 0.0008$\pm$0.00 0.0132$\pm$0.02
+AC 0.2208$\pm$0.13 0.2686$\pm$0.10 0.0935$\pm$0.11 0.0920$\pm$0.10 0.2430$\pm$0.08 0.2677$\pm$0.19 0.2402$\pm$0.08 0.1667$\pm$0.11 0.1266$\pm$0.12 0.2803$\pm$0.12 0.0009$\pm$0.00 0.3643$\pm$0.14
+VT 0.4575$\pm$0.03 0.4808$\pm$0.02 0.4808$\pm$0.02 0.4948$\pm$0.00 0.4673$\pm$0.02 0.4541$\pm$0.16 0.4893$\pm$0.00 0.4893$\pm$0.00 0.4844$\pm$0.00 0.4747$\pm$0.00 0.0001$\pm$0.00 0.5322$\pm$0.00
+$\overline{\textit{AR}}$ 8.79 6.75 6.25 6.00 7.04 5.58 5.96 5.50 7.17 7.25 9.7083 2.00
+Figure 6: Results of the two-tailed BD tests w.r.t. the NMI performance of different clustering approaches shown in Table 7.
+A.6 Ablation Study Evaluated by ARI and NMI
+The ablation study compares different ablated variants of OCL from the algorithm and data perspectives using the ARI and NMI indices, and the results are demonstrated in Table 8 - Table 11.
+The observations are as follows.
+From the algorithm aspect, it can be observed that the performance of OCL is superior to all variants on both ARI and NMI indices.
+More specifically, OCL outperforms OCL-I significantly on 10 out of 12 datasets both on ARI and NMI indices, and its performance is similar to OCL-I on the other 2 datasets.
+This indicates that using the probability-based order distance structure can effectively utilize the intra-cluster distribution information for more reasonable clustering-oriented distance measurement.
+Furthermore, the performance of OCL-I is also significantly better than OCL-II on 10 out of 12 datasets, which verifies the effectiveness of the order learning mechanism.
+In addition, compared to the OCL-III variant that completely disregards order information, OCL-II has better performance on all 12 datasets, which proves the effectiveness of considering orders for clustering.
+Table 8: ARI performance of different ablated versions of OCL on all the 12 datasets. “$\overline{\textit{AR}}$” column reports the average performance rankings.
+OCL Components Mark Datasets $\overline{\textit{AR}}$
 OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
-$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9830$\uparrow$   0.3573$\uparrow$   0.6222$\uparrow$   0.7458$\uparrow$   0.6313   0.4326$\uparrow$   0.7792$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   0.5785$\uparrow$   0.8206$\uparrow$   0.8943$\uparrow$   1.17$\uparrow$
-$\checkmark$ $\checkmark$ OCLI   0.9149$\uparrow$   0.3364$\uparrow$   0.6085$\uparrow$   0.7283$\uparrow$   0.6500$\uparrow$   0.3765$\uparrow$   0.7525$\uparrow$   0.5643$\uparrow$   0.4872$\uparrow$   0.5784$\uparrow$   0.7546$\uparrow$   0.8942$\uparrow$   2.00$\uparrow$
-$\checkmark$ OCLII   0.8681$\uparrow$   0.3359$\uparrow$   0.6069$\uparrow$ 0.
-7125$\uparrow$   0.6388$\uparrow$ 0.3712   0.7000$\uparrow$   0.5636$\uparrow$   0.4453$\uparrow$   0.5762$\uparrow$   0.7545$\uparrow$   0.8828$\uparrow$   2.92$\uparrow$
-OCLIII 0.8468 0.3208 0.5392 0.6525 0.5463 0.3758 0.6624 0.5472 0.4277 0.5569 0.7281 0.8625 3.92
-From the data aspect, to validate our hypothesis that correct order information is crucial to categorical data clustering, we conduct ablation experiments on datasets containing both nominal and ordinal attributes.
-We compare the proposed OCL, which learns orders for both the nominal and ordinal attributes, to two different variants.
-The first one is LNRO (Learn Nominal Reserve Ordinal), which retains the semantic order information of the original ordinal attribute and only learns orders for the nominal attributes.
-The other one is RNRO (Reserve Nominal Reserve Ordinal), which discards the order learning process for both nominal and ordinal attributes.
-Table 6 compares the clustering performance of the three variants.
-It can be observed that on five out of the six datasets, the performance of LNRO is superior to RNRO.
-This verifies our hypothesis that the nominal attributes also contain potential order information that is effective for clustering.
-Significantly, we can find that the OCL method outperforms LNRO on five out of the six datasets.
-This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering than the original semantic orders.
-Table 6. CA performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
-Learn Order for Variants Datasets $\overline{\textit{AR}}$
+$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9620$\uparrow$   0.1015$\uparrow$   0.0524$\uparrow$   0.2590$\uparrow$ 0.0627   0.0368$\uparrow$   0.7536$\uparrow$   0.0799$\uparrow$   0.1552$\uparrow$   0.0226$\uparrow$   0.4313$\uparrow$ 0.6207   1.25$\uparrow$
+$\checkmark$ $\checkmark$ I   0.8718$\uparrow$ 0.0592   0.0388$\uparrow$   0.2407$\uparrow$   0.0780$\uparrow$   0.0079$\uparrow$   0.7160$\uparrow$   0.0248$\uparrow$   0.1054$\uparrow$ 0.0132   0.2880$\uparrow$   0.6208$\uparrow$   2.08$\uparrow$
+$\checkmark$ II   0.8371$\uparrow$   0.0614$\uparrow$   0.0384$\uparrow$   0.2295$\uparrow$   0.0670$\uparrow$   -0.0003$\uparrow$   0.6482$\uparrow$   0.0247$\uparrow$   0.0953$\uparrow$   0.0205$\uparrow$   0.2879$\uparrow$   0.5850$\uparrow$   2.67$\uparrow$
+III 0.7638 0.0376 0.0022 0.1077 -0.0023 -0.0081 0.5787 0.0046 0.0843 0.0170 0.2596 0.5247 3.92
+Table 9: NMI performance of different ablated versions of OCL on all the 12 datasets. “$\overline{\textit{AR}}$” column reports the average performance rankings.
+OCL Components Mark Datasets $\overline{\textit{AR}}$
+OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
+$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9757$\uparrow$   0.1492$\uparrow$   0.0955$\uparrow$   0.2732$\uparrow$ 0.0808   0.0456$\uparrow$   0.8332$\uparrow$   0.0281$\uparrow$   0.1919$\uparrow$   0.0132$\uparrow$   0.3643$\uparrow$ 0.5322   1.25$\uparrow$
+$\checkmark$ $\checkmark$ I   0.9262$\uparrow$ 0.1086   0.0919$\uparrow$   0.2567$\uparrow$   0.0975$\uparrow$   0.0236$\uparrow$   0.8111$\uparrow$   0.0112$\uparrow$   0.1572$\uparrow$ 0.0062   0.2238$\uparrow$   0.5422$\uparrow$   2.08$\uparrow$
+$\checkmark$ II   0.9113$\uparrow$   0.1122$\uparrow$   0.0744$\uparrow$   0.2462$\uparrow$   0.0816$\uparrow$   0.0153$\uparrow$   0.7965$\uparrow$   0.0110$\uparrow$   0.1410$\uparrow$   0.0106$\uparrow$   0.2237$\uparrow$   0.5098$\uparrow$   2.67$\uparrow$
+III 0.8460 0.0458 0.0147 0.1352 0.0157 0.0088 0.7172 0.0044 0.1309 0.0105 0.2208 0.4575 3.92
+From the data aspect, the results can be observed that, on five out of six datasets, the ARI performance of LNRO is superior to RNRO.
+This verifies our hypothesis that the nominal attribute values also contain potential order information that is effective for clustering analysis.
+Although for the NMI index, LNRO outperforms RNRO on four out of six datasets, the performance gap between LNRO and RNRO is tiny, which can still verify the effectiveness of learning orders for nominal attributes in general.
+Significantly, we can find that the OCL method outperforms LNRO on five out of six datasets w.r.t.
+both ARI and NMI.
+This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering analysis than the ones directly derived from the semantics of attribute values.
+Table 10: ARI performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
+Learn Order For Mark Datasets $\overline{\textit{AR}}$
 OA NA NS AP CS HR BC LG
-OCL   0.3573$\uparrow$   0.6222$\uparrow$ 0.6313   0.4326$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   1.17$\uparrow$
-LNRO 0.3281   0.6074$\uparrow$   0.6350$\uparrow$   0.3985$\uparrow$   0.5476$\uparrow$   0.4818$\uparrow$   2.00$\uparrow$
-RNRO 0.3308 0.5984 0.6225 0.3902 0.5276 0.4493 2.83
-4.4. Convergence and Efficiency Evaluation
+OCL   0.1015$\uparrow$   0.0524$\uparrow$ 0.0627   0.0368$\uparrow$   0.0799$\uparrow$   0.1552$\uparrow$   1.17$\uparrow$
+LNRO 0.0705   0.0385$\uparrow$   0.0647$\uparrow$   0.0131$\uparrow$   0.0156$\uparrow$   0.1122$\uparrow$   2.00$\uparrow$
+RNRO 0.0734 0.0315 0.0545 0.0066 0.0062 0.1076 2.83
+Table 11: NMI performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
+Learn Order For Mark Datasets $\overline{\textit{AR}}$
+OA NA NS AP CS HR BC LG
+OCL   0.1492$\uparrow$   0.0955$\uparrow$ 0.0808   0.0456$\uparrow$   0.0281$\uparrow$   0.1919$\uparrow$   1.17$\uparrow$
+LNRO 0.1068   0.0791$\uparrow$   0.0840$\uparrow$   0.0230$\uparrow$   0.0083$\uparrow$ 0.1500   2.17$\uparrow$
+RNRO 0.1098 0.0619 0.0726 0.0189 0.0047 0.1610 2.67
+A.7 Efficiency Evaluation on Nursery School (NS) Dataset
+To make a more comprehensive evaluation of the execution efficiency of OCL, we also sample the real benchmark NS dataset with sampling rates from 50% to 500% with a step size of 50%, to simultaneously investigate the execution time in terms of $n$ and the clustering accuracy of OCL.
+Please note that the labels of samples are also sampled, and are only used to evaluate clustering accuracy.
+Since sampling attributes are meaningless for evaluating clustering accuracy, only the results w.r.t.
+number of data samples $n$ are reported here.
+The results are reported in Figure 7.
+Overall, the execution time of OCL is better than UDMC and AMPHM on different-sized NS datasets.
+It can also be observed that the clustering accuracy of the proposed OCL method is stable under different sample scale, and is always higher than the compared methods.
+Additionally, as the number of samples increases, the execution time of OCL increases linearly, which is consistent with our time complexity analysis in Section 3.5.
+In summary, OCL is effective and efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
+Figure 7: Clustering accuracy and execution time on NS dataset under different sampling rates $\{50\%,100\%,150\%,...,500\%\}$ with a step size of 50%.
+A.8 Case Study on Hayes-Roth (HR) Dataset
+It is worth noting that the proposed order learning method not only improves the clustering performance but also mines the hidden order information when facing datasets with unclear semantic information.
+This case study investigates whether the value order learned by the proposed method is interpretable and consistent with human subjective understanding.
+The Hayes-Roth dataset with one nominal and three ordinal attributes records personal information, such as the nominal “Hobby”, ordinal “Education level”, “Age”, and “Marital status”.
+The true labels of this dataset divide the samples into three groups of club members for psychological evaluation.
+Original dataset information is visualized in Figures 8 (a).
+It can be seen from Figure 8 (b) that our method can correctly learn the semantic order of the values from attributes “Education level” and “Age”, as the learned orders are consistent with the original semantic orders.
+Moreover, for the nominal attribute “Hobby”, although there is no original obvious semantic order, we show below that the learned order is very meaningful in data analysis and knowledge acquisition.
+First, we can find that the learned order distance between “chess” and “stamp collection” hobbies is closer, while the distance between “stamp collection” and “sport” is larger.
+This is consistent with a cognitive intuition that “chess” is a quieter sport, so it ranks between the quieter hobby “stamp collection” and its sibling “sport”.
+This indicates that OCL can learn latent semantic orders for nominal attributes, which is an important characteristic for categorical data understanding.
+As shown in the clustering results obtained by OCL in Figure 8 (c), it can be seen that there are more cases of people who enjoy playing chess and collecting stamps appearing in the same cluster, which is consistent with human intuition and the learned order of “Hobby” attribute.
+This further confirms the prospects of our OCL method in discovering the implicit order information of unlabeled data with no obvious semantic order.
+For the “Marital status” attribute, the order learned by OCL is inconsistent with the original semantic order as demonstrated in Figure 8 (b).
+The order learned by OCL indicates that the possible values semantically ranked first and third in “Marital status” should be more similar to better serve clustering.
+From the true clusters given by the dataset shown in Figure 8 (d), it can be found that more of the first and third values of the “Marital status” attribute simultaneously appear in the same clusters, which confirms the reasonableness of the learned order.
+This excellent characteristic allows the users to better understand the cluster distribution of data by using the proposed OCL method.
+[…]
+A.9 Convergence Curves of All 12 Datasets
 To evaluate the convergence of OCL, we execute it on all 12 datasets and present the convergence curves in Figure 9.
@@
 This reflects that OCL converges efficiently with guarantee, which conforms to the analysis in Section 3.5.
-Figure 3. Convergence curve of OCL on all 12 datasets.
-Furthermore, to evaluate the execution efficiency of OCL, we construct large synthetic datasets with different numbers of samples and attributes to investigate the execution time of OCL in comparison with other counterparts.
-The results reported in Figure 4 show that the execution time of OCL is better than UDMC and DLC on large-scale datasets, and better than H2H, UDMC, and CBDM on high-dimensional datasets.
-Additionally, as the number of samples or attributes increases, the execution time of OCL increases linearly, which is consistent with the time complexity analysis in Section 3.5.
-In summary, OCL is efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
-Figure 4. Execution time on synthetic datasets with different numbers of samples $n$ and attributes $d$.
-4.5. Visualization of Clustering Effect
-To demonstrate the cluster discrimination ability of the OCL method, we utilize the distance between attribute values learned by OCL, CBDM, and ADC respectively to encode the attributes of the VT dataset.
-Subsequently, the t-SNE (Maaten and Hinton, 2008) is introduced to process the encoded data into two-dimensional space and visualize it by marking the samples with real labels in different colors.
-It can be seen from Figure 5 that OCL has significantly better cluster discrimination ability than CBDM and ADC.
-(a) CBDM @@@@ VT
-(b) ADC @@@@ VT
-(c) OCL @@@@ VT
-Figure 5. t-SNE visualization of the VT dataset represented by CBDM, ADC, and OCL.
-5. Concluding Remarks
-This paper introduces a new finding that knowing the correct order relation of attribute values is critical to cluster analysis of categorical data.
-Accordingly, we propose the OCL paradigm that allows joint learning of data partitions and order relation of attribute values for accurate clustering.
-The proposed learning mechanism can infer a more appropriate order among categorical attribute values based on the current clusters, thus yielding more compact clusters based on the updated order.
-The data partition and order learning are iteratively executed to approach the clustering objective, i.e., maximizing the overall similarity of intra-cluster samples.
-As a result, more accurate clustering results and insightful orders can be obtained simultaneously.
-OCL converges quickly with guarantee, and extensive experiments conducted on 12 real benchmark datasets fully demonstrate its superiority in comparison with state-of-the-art clustering methods.
-A case study has also been provided to show the rationality of the learned order in understanding unlabeled categorical data.
-The proposed OCL method is also not exempt from limitations.
-Although it demonstrates the potential to enhance the performance of mixed data clustering, the deep connection between the treatment of categorical and numerical attributes remains to be explored.
-In addition, it would also be promising to extend OCL to more complex datasets with non-static distribution, an unknown number of clusters, and extremely imbalanced clusters.
+Figure 9: Convergence curve of OCL on all 12 datasets.
+A.10 Clustering Performance on Mixed Datasets
+To demonstrate the potential of OCL in boosting the clustering of mixed data comprising both numerical and categorical attributes, we compare the original $k$-prototypes (Huang, 1997), HARR (Zhang et al., 2022a), and ADC (Zhang & Cheung, 2023) methods proposed for mixed data clustering, with their corresponding variants enhanced by the orders learned by OCL.
+More specifically, for $k$-prototypes (KPT), we first encode the categorical attribute values using the orders learned by the proposed OCL method and then treat the dataset as a pure numerical dataset for clustering using the $k$-means (KMS) algorithm (Ball & Hall, 1967).
+Actually, KPT enhanced by OCL becomes KMS with the categorical attributes encoded into numerical ones by OCL, and thus we call the enhanced version KMS+OCL.
+For HARR and ADC, since they are originally competent in processing ordinal attributes, we assign the orders learned by OCL to the categorical attributes, and let HARR and ADC treat all the categorical attributes as ordinal attributes.
+Accordingly, their enhanced versions are named HARR+OCL and ADC+OCL, respectively.
+Table 12 demonstrates the performance comparison of the above-mentioned mixed data clustering approaches.
+It can be observed that OCL considerably enhances the clustering performance of the three approaches in terms of all three commonly used validity indices on AP and AC datasets.
+This significantly indicates that the learned orders can effectively improve the clustering performance on mixed data.
+It can also be observed that the improvement on the AP dataset is more obvious.
+The reason would be that AP contains both nominal and ordinal attributes as shown in Table 6.
+This situation allows OCL to leverage its superiority in both learning orders for nominal attributes and potential corrections on the semantic orders of ordinal attributes.
+In addition, a noteworthy special case can be seen on the DS dataset.
+Since DS has only two possible values on each of its categorical attributes, OCL cannot further optimize the orders through learning, because the order distance between the only two possible values is naturally fixed.
+This is why the clustering performance on the DS dataset remains unchanged after using the orders learned by OCL.
+This case brings us a hint that if the original semantic order of an ordinal attribute is already optimal, OCL will not have an impact on boosting the clustering performance.
+In short, the proposed OCL is promising in being applied to more complex mixed data clustering where the attribute values originally have no order or with implicit order.
+Table 12: Clustering performance of mixed data clustering approaches and their OCL-enhanced versions i.e., adopting the distance metric learned by OCL on AP, AC, and DS datasets.
+The green arrow indicates that an improvement has been achieved by the corresponding OCL-enhanced version.
+Datasets Methods Metric
+CA ARI NMI
+AP KMS+OCL 0.5550$\uparrow$ 0.0104$\uparrow$ 0.0106$\uparrow$
+KPT 0.5323 0.0015 0.0051
+HARR+OCL 0.6042$\uparrow$ 0.0373$\uparrow$ 0.0593$\uparrow$
+HARR 0.5688 0.0164 0.0371
+ADC+OCL 0.5677$\uparrow$ 0.0144$\uparrow$ 0.0126$\uparrow$
+ADC 0.5487 0.0049 0.0070
+AC KMS+OCL 0.7975$\uparrow$ 0.3537$\uparrow$ 0.2738$\uparrow$
+KPT 0.7425 0.2794 0.2154
+HARR+OCL 0.8232$\uparrow$ 0.4169$\uparrow$ 0.3225$\uparrow$
+HARR 0.8213 0.4121 0.3184
+ADC+OCL 0.6480$\uparrow$ 0.0839$\uparrow$ 0.0644$\uparrow$
+ADC 0.6445 0.0797 0.0607
+DS KMS+OCL 0.6842 0.1684 0.1888
+KPT 0.6842 0.1684 0.1888
+HARR+OCL 0.8117 0.4364 0.4277
+HARR 0.8117 0.4364 0.4277
+ADC+OCL 0.8667 0.5336 0.4244
+ADC 0.8667 0.5336 0.4244
+A.11 Clustering Performance on Numerical Datasets
+Experimental Settings
+To demonstrate the potential of our OCL for numerical data clustering, we compare OCL on two numerical datasets obtained from the UCI machine learning repository (http://archive.ics.uci.edu/ml).
+The dataset statistics and comparative results are shown in Table 13.
+To implement OCL on these two numerical datasets, we adopt uniform discretization for the numerical attributes.
+That is, the values of each numerical attribute are divided into ten intervals of equal length, and these ten intervals are regarded as ten possible values with an order relationship.
+We implement OCL on the discretized numerical datasets to learn new orders for the intervals.
+To illustrate the effectiveness of order learning, we also preserve the original order by implementing OCL without order learning (OCL w.o.
+OL).
+More specifically, OCL w.o.
+OL is a version of OCL with fixed original orders.
+It is equivalent to executing only the inner loop of Algorithm 1 in our paper to learn clusters.
+The outer loop does not make any adjustments to the order to ensure that it is fixed.
+The conventional K-Means clustering algorithm that utilizes the original order of numerical data by default is also compared.
+All the results are averaged on ten runs of the compared methods.
+Clustering Performance
+The comparative results are shown in Table 13.
+It can be seen that the order learned by OCL can surpass the original order preserved by the K-Means and OCL w.o.
+OL in general.
+This is because the order of a numerical attribute is not necessarily linearly related to the true clusters.
+For example, people with medium incomes may have happier family lives than those with very high or meager incomes.
+Therefore, the original order of the “Income” attribute may not be beneficial for forming two clusters of people with happy and unhappy family lives.
+In this scenario, the proper income order should be something like: medium $>$ high $>$ low, or reverse.
+The order learned by OCL can reflect such non-linear relationships among values, and thus yields more accurate clustering results in the corresponding situations.
+Table 13: Clustering performance on two numerical datasets, $d$, $n$, and $k^{*}$ are the numbers of attributes, samples, and true clusters, respectively.
+The best results on each dataset are highlighted in bold.
+No. Dataset $d$ $n$ $k^{*}$ Metric K-Means OCL w.o. OL OCL
+1 Shuttle 9 43500 7 CA 0.4217$\pm$0.00 0.5358$\pm$0.03 0.6415$\pm$0.02
+ARI 0.1912$\pm$0.00 0.2234$\pm$0.02 0.2640$\pm$0.01
+NMI 0.4349$\pm$0.00 0.3566$\pm$0.01 0.3115$\pm$0.01
+2 Avail 10 20867 12 CA 0.2239$\pm$0.00 0.2487$\pm$0.00 0.3212$\pm$0.00
+ARI 0.0201$\pm$0.00 0.0301$\pm$0.00 0.0323$\pm$0.00
+NMI 0.0854$\pm$0.00 0.0920$\pm$0.00 0.0953$\pm$0.00
+Detailed Observations
+Three detailed observations are provided in the following: (1) OCL performs better than the two counterparts with fixed original orders (i.e., OCL w.o.
+OL and K-Means) in general.
+This proves the effectiveness of the orders learned by OCL on the two numerical datasets.
+This also verifies our guess that order learning can be extended to numerical data clustering.
+By carefully observing the data attribute values and cluster distributions, we found that the nonlinear relationship between attributes and the label attribute is the key factor causing this phenomenon.
+However, the main challenge lies in that the labels of the data for clustering are all categorical.
+How to reflect the implicit relationship between heterogeneous numerical attribute and categorical label attribute is the key to deploying order learning for numerical data clustering.
+(2) Both K-Means and OCL w.o.
+OL use the original order of numerical attributes, but the latter performs better in general.
+This is because the latter OCL w.o.
+OL uses a more advanced learnable distance, which adaptively adjusts the distances between attribute values during clustering.
+In contrast, K-Means uses the Euclidean distance, which is given in advance and thus has no connection with the clustering task.
+(3) The NMI performance of K-Means is obviously higher than that of the two OCL variants (i.e., the full version of OCL and OCL w.o.
+OL) on the Shuttle dataset.
+By observing the distributions of attribute values and clusters in detail, we found that the distribution of the shuttle dataset is extremely imbalanced.
+The values of many attributes are concentrated in a very small interval, which is very unfavorable for the equal interval discretization adopted by the two OCL variants.
+Therefore, in comparison, the Euclidean distance used by K-Means can retain more discrimination information between samples, so the corresponding clustering performance is better in terms of the NMI metric.
+On the other hand, the cluster distribution of this dataset is also extremely imbalanced (the imbalance ratio, i.e.
+the number of samples in the largest true cluster divided by the number of samples in the smallest true cluster, is as high as 5684.7!).
+The extreme imbalance will lead to unstable verification results of the ARI and NMI metrics.
+This explains why the compared methods exhibit very different performance positions on ARI and NMI.
+However, since the CA metric first performs the optimal mapping from the obtained clusters to the label clusters and then evaluates the performance, CA will be relatively reliable in the extremely imbalanced case and thus still verifies the effectiveness of order learning of numerical data.
+Table 14: Dataset statistics of the six larger-scale datasets.
+$d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, $n$, and $k^{*}$ are the numbers of nominal attributes, ordinal attributes, samples, and true clusters, respectively.
+The full name of the “Coupon” dataset is “In-Vehicle Coupon Recommendation”.
+No. Dataset Abbrev. $d^{\text{\{nom\}}}$ $d^{\text{\{ord\}}}$ $n$ $k^{*}$
+1 Obesity Levels OB 6 2 2111 7
+2 Auction Verifications AV 0 6 2043 2
+3 Chess CC 36 0 3195 2
+4 Coupon CR 7 4 12684 2
+5 Bank Marking BM 6 3 45211 2
+6 Connect-4 CT4 42 0 67557 3
+Table 15: Clustering performance evaluated by CA, NMI, and ARI on the six larger-scale datasets.
+The best and second-best results on each dataset are highlighted in bold and underlined, respectively.
+The “$\overline{\textit{AR}}$” row reports the average performance ranks of each method in all 18 comparisons.
+Data Metric KMD LSM JDM CBDM UDMC DLC HARR HDC ADC MCDC AMPHM OCL
+CT4 CA 0.4000$\pm$0.02 0.3609$\pm$0.01 0.3696$\pm$0.02 0.3651$\pm$0.02 0.3819$\pm$0.03 0.3618$\pm$0.01 0.3924$\pm$0.03 0.3879$\pm$0.03 0.3914$\pm$0.03 0.3514$\pm$0.01 0.3824$\pm$0.00 0.4121$\pm$0.02
+ARI -0.0027$\pm$0.00 0.0010$\pm$0.00 0.0001$\pm$0.00 0.0004$\pm$0.00 0.0013$\pm$0.00 0.0012$\pm$0.00 $0.0$$0.0$0005$\pm$0.00 0.0002$\pm$0.00 0.0005$\pm$0.00 0.0150$\pm$0.01
+NMI 0.0028$\pm$0.00 0.0029$\pm$0.00 0.0017$\pm$0.00 0.0015$\pm$0.00 0.0036$\pm$0.00 0.0013$\pm$0.00 0.0011$\pm$0.00 0.0016$\pm$0.00 0.0016$\pm$0.00 0.0053$\pm$0.00 0.0023$\pm$0.00 0.0128$\pm$0.01
+OB CA 0.3715$\pm$0.03 0.3697$\pm$0.02 0.3594$\pm$0.03 0.3673$\pm$0.03 0.3720$\pm$0.02 0.3724$\pm$0.04 0.3797$\pm$0.02 0.3617$\pm$0.04 0.3765$\pm$0.03 0.3717$\pm$0.02 0.3512$\pm$0.00 0.3935$\pm$0.01
+ARI 0.1527$\pm$0.03 0.1588$\pm$0.02 0.1347$\pm$0.04 0.1522$\pm$0.04 0.1580$\pm$0.02 0.1548$\pm$0.04 0.1727$\pm$0.02 0.1417$\pm$0.05 0.1472$\pm$0.04 0.1527$\pm$0.04 0.1672$\pm$0.00 0.1696$\pm$0.01
+NMI 0.2256$\pm$0.03 0.2330$\pm$0.02 0.2081$\pm$0.04 0.2262$\pm$0.03 0.2342$\pm$0.02 0.2391$\pm$0.03 0.2644$\pm$0.02 0.2261$\pm$0.04 0.2233$\pm$0.03 0.2466$\pm$0.03 0.2333$\pm$0.00 0.2633$\pm$0.01
+AV CA 0.6251$\pm$0.08 0.6035$\pm$0.10 0.6291$\pm$0.12 0.6695$\pm$0.10 0.6065$\pm$0.10 0.6150$\pm$0.06 0.6441$\pm$0.14 0.6364$\pm$0.10 0.6325$\pm$0.12 0.6114$\pm$0.10 0.6125$\pm$0.00 0.6637$\pm$0.07
+ARI 0.0073$\pm$0.02 0.0111$\pm$0.02 0.0136$\pm$0.03 0.0365$\pm$0.02 0.0123$\pm$0.03 0.0262$\pm$0.03 0.0214$\pm$0.03 0.0199$\pm$0.02 0.0173$\pm$0.02 0.0209$\pm$0.02 0.0166$\pm$0.00 0.0393$\pm$0.02
+NMI 0.0021$\pm$0.00 0.0024$\pm$0.00 0.0024$\pm$0.00 0.0099$\pm$0.01 0.0027$\pm$0.00 0.0093$\pm$0.01 0.0075$\pm$0.01 0.0035$\pm$0.00 0.0034$\pm$0.00 0.0025$\pm$0.01 0.0044$\pm$0.00 0.0110$\pm$0.01
+BM CA 0.5730$\pm$0.06 0.5792$\pm$0.04 0.5959$\pm$0.07 0.6011$\pm$0.07 0.5709$\pm$0.03 0.5944$\pm$0.05 0.6064$\pm$0.12 0.5457$\pm$0.03 0.5543$\pm$0.04 0.5573$\pm$0.02 0.5039$\pm$0.00 0.6127$\pm$0.00
+ARI 0.0146$\pm$0.03 0.0107$\pm$0.02 0.0183$\pm$0.04 0.0054$\pm$0.05 0.0074$\pm$0.02 0.0178$\pm$0.02 0.0186$\pm$0.01 -0.0037$\pm$0.02 0.0035$\pm$0.02 -0.0007$\pm$0.02 0.0072$\pm$0.00 0.0233$\pm$0.00
+NMI 0.0211$\pm$0.00 0.0143$\pm$0.01 0.0135$\pm$0.01 0.0198$\pm$0.01 0.0179$\pm$0.01 0.0121$\pm$0.01 0.0155$\pm$0.01 0.0192$\pm$0.01 0.0204$\pm$0.01 0.0103$\pm$0.01 0.0177$\pm$0.00 0.0194$\pm$0.00
+CC CA 0.5562$\pm$0.04 0.5572$\pm$0.03 0.5632$\pm$0.03 0.5433$\pm$0.03 0.5587$\pm$0.02 0.5342$\pm$0.04 0.5248$\pm$0.02 0.5438$\pm$0.03 0.5395$\pm$0.04 0.5668$\pm$0.03 0.5012$\pm$0.00 0.5726$\pm$0.05
+ARI 0.0165$\pm$0.01 0.0166$\pm$0.01 0.0189$\pm$0.02 0.0103$\pm$0.01 0.0150$\pm$0.01 0.0099$\pm$0.02 0.0019$\pm$0.01 0.0108$\pm$0.01 0.0113$\pm$0.02 0.0177$\pm$0.01 0.0106$\pm$0.00 0.0285$\pm$0.02
+NMI 0.0150$\pm$0.01 0.0149$\pm$0.01 0.0149$\pm$0.01 0.0084$\pm$0.01 0.0119$\pm$0.01 0.0076$\pm$0.01 0.0023$\pm$0.01 0.0148$\pm$0.01 0.0141$\pm$0.02 0.0156$\pm$0.01 0.0144$\pm$0.00 0.0213$\pm$0.02
+CR CA 0.5209$\pm$0.02 0.5247$\pm$0.02 0.5185$\pm$0.02 0.5297$\pm$0.02 0.5245$\pm$0.01 0.5488$\pm$0.03 0.5541$\pm$0.00 0.5466$\pm$0.02 0.5504$\pm$0.02 0.5246$\pm$0.02 0.5510$\pm$0.00 0.5590$\pm$0.02
+ARI 0.0021$\pm$0.00 0.0033$\pm$0.00 0.0024$\pm$0.00 0.0039$\pm$0.00 0.0030$\pm$0.00 0.0120$\pm$0.01 0.0114$\pm$0.00 0.0101$\pm$0.01 0.0106$\pm$0.01 0.0096$\pm$0.00 0.0122$\pm$0.00 0.0145$\pm$0.01
+NMI 0.0017$\pm$0.00 0.0024$\pm$0.00 0.0020$\pm$0.00 0.0022$\pm$0.00 0.0022$\pm$0.00 0.0071$\pm$0.01 0.0115$\pm$0.00 0.0065$\pm$0.00 0.0057$\pm$0.00 0.0066$\pm$0.00 0.0048$\pm$0.00 0.0083$\pm$0.01
+$\overline{\textit{AR}}$ 7.58 7.28 8.06 6.92 6.97 6.39 4.94 7.53 7.08 6.75 7.11 1.39
+A.12 Additional Large-scale Datasets Results
+Six additional larger datasets in different fields have been chosen from the UCI machine learning repository (see http://archive.ics.uci.edu/ml) to further evaluate the clustering performance of the proposed OCL method.
+The statistics of the newly added datasets are summarized in Table 14.
+All datasets are pre-processed following the same settings in the submitted paper (see Section 4.1 and Appendix A.3).
+Clustering Performance
+The clustering performance of different methods is compared on the six additional datasets in Table 15.
+Consistent with the results in our paper, the proposed OCL outperforms the other methods in general.
+More detailed observations are provided below: (1) Average performance ranking of OCL is around 1.4 in 18 comparisons formed on 6 datasets under 3 metrics, while the most competitive rivals, i.e., DLC and HARR, rank around 6.4 and 4.9, respectively.
+The huge ranking gap (larger than 3) intuitively demonstrates the superiority of OCL.
+(2) All the experimental datasets are real public datasets from various fields, such as healthcare, gaming, recommendation systems, finance, etc.
+This configuration is sufficient to verify the superiority of OCL without bias.
+(3) The proposed OCL always ranks first or second in all comparisons (except for NMI performance on the BM dataset) and performs the best in 13 out of 18 comparisons, indicating that the advantage of OCL is very stable and extensive.